### v1

# Chain-of-Frames: Advancing Video Understanding in Multimodal LLMs via Frame-Aware Reasoning

Abstract
Recent work has shown that eliciting Large Language Models (LLMs) to generate reasoning traces in natural language before answering the user’s request can significantly improve their performance across tasks. This approach has been extended to multimodal LLMs, where the models can produce chain-of-thoughts (CoT) about the content of input images and videos. In this work, we propose to obtain video LLMs whose reasoning steps are grounded in, and explicitly refer to, the relevant video frames. For this, we first create CoF-Data, a large dataset of diverse questions, answers, and corresponding frame-grounded reasoning traces about both natural and synthetic videos, spanning various topics and tasks. Then, we fine-tune existing video LLMs on this chain-of-frames (CoF) data. Our approach is simple and self-contained, and, unlike existing approaches for video CoT, does not require auxiliary networks to select or caption relevant frames. We show that our models based on CoF are able to generate chain-of-thoughts that accurately refer to the key frames to answer the given question. This, in turn, leads to improved performance across multiple video understanding benchmarks, for example, surpassing leading video LLMs on Video-MME, MVBench, and VSI-Bench, and notably reducing the hallucination rate. Code available at github.com/SaraGhazanfari/CoF.
1 Introduction
Large Language Models (LLMs) are able to perform step-by-step reasoning, widely known as chain-of-thoughts (CoT) (Wei et al., 2022; Kojima et al., 2022). This capability has been implicitly integrated into state-of-the-art systems such as OpenAI’s o1/o3 models (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025), contributing to their remarkable performance and improving the interpretability of their internal functioning. CoT reasoning has been also extended to multimodal LLMs (Hu et al., 2024; Wu and Xie, 2023; Shao et al., 2024): this presents new challenges compared to language-only domains, as the models need to attend to inputs from different modalities, and reason about both their individual content and how they are connected (Awal et al., 2023; Kil et al., 2024; Sun et al., 2025).
Recent work has also begun to explore the integration of CoT into video understanding tasks, where the input to a multimodal LLM consists of a sequence of images (the frames of the video) along with a text prompt. This makes reasoning on videos particularly complex, as the model needs to capture the semantics of the text prompt, understand temporal and causal relationships between frames, and reason about the video in its entirety. Existing approaches rely on complex inference frameworks with architecture modifications (Fei et al., 2024) or auxiliary networks (Han et al., 2024) at evaluation time to integrate reasoning into video LLMs. This makes using these models both more computationally expensive and less general (as they are specialized to some tasks), and deviates from the natural CoT prompting successfully applied to standard LLMs. Another limitation of current methods is that collecting reasoning traces, needed for training, involves complex and costly procedures, which may limit how much the size of training datasets can be scaled in practice. For example, Wang et al. (2024b) iteratively refines the samples via annotations from both LLMs and human experts, Han et al. (2024) leverages several auxiliary models, Fei et al. (2024) uses additional spatial-temporal scene graph data. Finally, video LLMs do not currently provide explicit connections between particular segments of the video and their reasoning, which would better ground the chain-of-thought.
To remedy these limitations, in this work we propose chain-of-frames (CoF), a new frame-aware chain-of-thought reasoning approach for video LLMs that integrates temporal information directly into the CoT structure (see Fig. 1). This enables the model to identify and refer to the most relevant frames while answering questions, in contrast to prior works that treat frame selection and reasoning as separate stages. Chain-of-frames is a simple and natural adaptation of the CoT paradigm in NLP to video understanding that does not require the auxiliary networks or complex inference frameworks of existing methods. Moreover, we propose an efficient data generation pipeline that allows us to collect a large dataset of CoF examples, named CoF-Data. To achieve this, a key element consists of leveraging a synthetic video dataset (Yi et al., 2020) to extract a large and diverse set of reasoning traces at virtually no cost. Then, we fine-tune two recent open-source video LLMs, InternVL2.5-4B (Chen et al., 2024) and InternVL3-8B (Zhu et al., 2025), on our CoF-Data. In an extensive evaluation on five established benchmarks, we show that our CoF models significantly outperform the original ones, with and without naive CoT prompting (Wei et al., 2022) (see Fig. 1). Moreover, they are competitive with or better than state-of-the-art (SOTA) open- and closed-source models (Table 2). Notably, our CoF-InternVL3-8B achieves higher accuracy than the best model reported on Video-MME, MVBench, and VSI-Bench leaderboards, surpassing even proprietary models like GPT-4o and Gemini 1.5 Pro. These results also demonstrate that video LLMs can learn the ability to produce frame-aware reasoning from a limited number of (synthetic) samples and generalize the behavior to unseen tasks. Thus, our work provides a simple and inexpensive approach to adapt the CoT paradigm for video LLMs that improves both performance and interpretability.
(a) A CoF reasoning trace generated by our CoF-4B model.
(b) CoF models vs the baseline models.
Figure 1: (a) Chain-of-frames reasoning generated by our CoF-InternVL2.5-4B model: it includes the key frames to answer the question (from Video-MME). (b) Comparison of accuracy across multiple video understanding benchmarks between baseline models (InternVL2.5-4B and InternVL3-8B) and their CoF-enhanced counterparts: our models consistently outperform the baselines.
Figure 2: CoF-Data. We show examples of our training data with chain-of-frames reasoning, including video, question, answer and reasoning trace. We include samples from CoF-Datareal (real video, top row) and CoF-Datasynth (synthetic video, bottom), created as described in Sec. 3.3.
2 Related Work
Multimodal LLMs for videos. Multimodal Large Language Models (Chen et al., 2023b; Li et al., 2024b; Ghazanfari et al., 2024) have made substantial progress in integrating visual and textual modalities, enabling them to perform complex reasoning and achieve deep understanding across various data types, including videos (Chen et al., 2023b; Xue et al., 2024; Li et al., 2023a; Zhang et al., 2024b, c; Li et al., 2024a; Bai et al., 2024; Zhu et al., 2025). Among recent advancements, InternVL2.5 (Chen et al., 2024), InternVL3 (Zhu et al., 2025), LLaVA-NeXT-Video (Zhang et al., 2024c), and Qwen2-VL (Wang et al., 2024a) stand out for their strong video understanding capabilities. Both InternVL and LLaVA-NeXT-Video process videos in an image-text interleaved format, aligning sequences of video frames with language to form a unified multimodal stream. In addition to these open-source models, closed-source systems such as GPT-4o (Achiam et al., 2023) and Gemini-1.5 (Gemini Team et al., 2024) have also demonstrated impressive multimodal capabilities, although details of their architecture and training remain proprietary.
Chain-of-Thoughts for videos. Recent research on chain-of-thoughts for video understanding broadly falls into two categories: methods that provide explicit textual reasoning explanations (Wang et al., 2024b), and those that focus on identifying relevant frames to facilitate the process of generating the response (Han et al., 2024; Hu et al., 2025b, a). In the first category, VideoCoT (Wang et al., 2024b) introduces an active annotation tool to generate reasoning explanations, thereby encouraging models to explicitly reason on visual content. In the second category, several approaches emphasize frame selection as a pre-processing step to enhance reasoning efficiency. M-LLM (Hu et al., 2025b) proposes leveraging multimodal LLMs to identify the most relevant frames corresponding to the query. Likewise, Chain-of-Shot (Hu et al., 2025a) introduces a prompting strategy specifically tailored for understanding long-form videos by selecting key frames. VideoEspresso (Han et al., 2024) combines the ideas of core frame selection with fine-grained reasoning annotations, creating a large-scale dataset that supports more efficient and focused video reasoning. Finally, Fei et al. (2024) introduces Video-of-Thought, a complex five-step pipeline to generate spatial-temporal scene graphs to answer multiple-choice questions about videos.
3 Chain-of-Frames: Reasoning on Videos via Frame References
3.1 Limitations of reasoning on videos
Multimodal LLMs (Zhang et al., 2024c; Chen et al., 2024; Bai et al., 2024) process videos as a sequence of images (frames), which are encoded as image tokens by a vision encoder, then concatenated with the user prompt in natural language, and finally passed to a language model. While the base language models may be trained to produce reasoning traces, these are not specific for reasoning on videos.
To encourage chain-of-thought output in video LLMs, models must be fine-tuned on video-grounded reasoning traces. Wang et al. (2024b) introduces VideoCoT, which used reasoning traces generated by LLMs and refined by human experts to describe video events. However, the traces lack explicit temporal grounding, that is, individual reasoning steps are not clearly aligned with the corresponding video frames. Additionally, the reliance on human and LLM annotations makes data generation expensive, limiting the dataset to only 11k samples. A different strategy is proposed by Han et al. (2024), who use multiple auxiliary models to generate CoT data. Lightweight multimodal LLMs select the core frames, GPT-4o (Hurst et al., 2024) identifies key elements in each core frame and their relevance to the query, GroundingDINO (Liu et al., 2024) provides spatial annotations (e.g., bounding boxes), and the BGE-M3 retriever (Chen et al., 2023a) produces the temporal annotations. While effective during training, this approach cannot be deployed at inference time given its high complexity and cost (only the frame selection step, which depends on auxiliary LLMs, is kept). Alternatively, Video-of-Thought (Fei et al., 2024) introduces a complex five-step inference pipeline, which includes generating the spatial-temporal scene graphs for key frames and is specialized for multiple-choice questions.
In summary, existing approaches to video reasoning with LLMs face three main limitations: (i) expensive training data generation, (ii) complex inference, possibly involving auxiliary models, and (iii) lack of explicit temporal grounding in the reasoning process.
3.2 Chain-of-Frames
To address these limitations, we propose Chain-of-Frames (CoF), a simple yet effective approach introducing temporal grounding into the reasoning process. CoF consists of reasoning traces with explicit references to frames relevant to answering the given query. Concretely, we use the position of the frame in the video (e.g., “Frame 1”, “Frame 2”, …) as an identifier. Unlike timestamps, this representation is agnostic to video duration and sample frequency, making it more consistent across diverse video data and potentially easier to learn. Examples are shown in Fig. 1 and Fig. 2, with additional illustrations provided in App. C. This approach has several benefits
  1. 1.
    Data availability: as we show in Sec. 3.3, training data can be efficiently generated from existing annotated real videos. Additionally, CoF traces can be created from synthetic videos at virtually no cost. This low-cost data generation contrasts with the more complex pipelines of Wang et al. (2024b); Han et al. (2024).
  2. 2.
    Simplicity: the reasoning traces are completely in natural language (without the need for complex format, bounding boxes or scene graphs as for Fei et al. (2024); Han et al. (2024)). This approach is a natural extension of standard CoT method used in NLP tasks and does not require specific inference pipelines.
  3. 3.
    Temporal grounding: by explicitly referencing the frames relevant to the task, CoF increases the connection between video and reasoning. Thus, our method can be seen as a temporally grounded extension of the typical CoT of Wang et al. (2024b).
  4. 4.
    Interpretability: the reasoning traces generated at inference time include frame references (Fig. 1) and provide direct insights into how the LLM’s decisions are obtained.
In the next section, we detail how we generate chain-of-frames from video datasets. These traces constitute the training data for fine-tuning video LLMs capable of producing CoF-based reasoning.
Figure 3: Overview of our two-step pipeline for generating CoF-Data. Step 1 adjusts the frame IDs while preserving frame-caption alignment. Step 2 utilizes raw annotations to generate CoF triplets (question, frame-aware reasoning trace, answer). For this, we leverage Llama3.1-8B a manual template with the synthetic videos from Clevrer.
3.3 Chain-of-Frames training data collection (CoF-Data)
We construct chain-of-frames traces from both real and synthetic videos. For real videos, we use the training split of the VideoEspresso dataset (Han et al., 2024), which features videos from a wide range of sources and includes descriptions of key frames for each. These diverse frame-level annotations provide a starting point for extracting CoF traces. To complement this, we use the training split of the Clevrer dataset (Yi et al., 2020), which contains synthetic videos of simple 3D objects interacting within a controlled environment, along with rich annotations. Synthetic data offers two key advantages: it overcomes the cost of generating training samples from real videos and mitigates the scarcity of sufficiently annotated real videos. We next describe the main steps of our data generation pipeline (illustrated in Fig. 3), and introduce the resulting dataset of chain-of-frames, named CoF-Data. Additional details and examples are provided in App. A.
Frame ID alignment. The original annotations include frame IDs, but, due to context length limitations of video LLMs, we downsample the videos while preserving the frame-annotation alignment. We first map each frame to its timestamp, and clip the video to the maximum duration allowed by the model (e.g., 30 seconds in our experiments), ensuring the segment includes all frames for which captions are available. We then re-calibrate the frame IDs to reflect their new positions within the clipped video.
CoF from real videos (CoF-Datareal). VideoEspresso provides captions for key frames. After aligning frame IDs, we obtain data in the format shown in Fig. 3 (raw annotations). From these annotations, we generate triplets of questions, answers, and reasoning traces with frame references by prompting an LLM using the raw annotations as input. For this, we use Llama3.1-8B-Instruct (Meta AI, 2024) (the full prompt is provided in App. A). This process yields multiple questions per video, often covering diverse parts of the video and referencing different sets of frames. An example from CoF-Datareal is illustrated in Fig. 3, and a complete training sample including the video is in Fig. 2.
CoF from synthetic videos (CoF-Datasynth). In each frame of Clevrer, every object is annotated with both fixed properties (shape, material, color) and situational attributes (e.g., velocity, location), see Fig. 3. We use these rich attributes to generate three categories of quantitative questions (object count, appearance order, and relative distance) which complement the semantic questions obtained for real video. Notably, we can generate both questions, answers, and chain-of-frames using fixed manual templates (see App. A), since all the necessary information can be directly deduced from the object-specific raw annotations. This eliminates the need for using an LLM, significantly reducing generation cost and enabling easy scaling of the dataset size. An example of an appearance order question is shown in Fig. 2 (examples from other categories are shown in App. C).
Figure 4: Distribution of frame references in the Chain-of-Frames training data. The left pie chart illustrates the distribution for CoF-Datareal, having fewer frames per reasoning trace, whereas CoF-DATA-synth demonstrates a more balanced frame distribution due to controlled synthetic video generation. The right pie chart shows the overall distribution for the CoF-Data.
Final dataset (CoF-Data). From the generated chain-of-frames, we filter out samples where frames are referred in the question (as this does not happen at test time). Moreover, we reduce the number of samples with no frame references in the reasoning trace to give higher weight to more complex examples of reasoning. We nevertheless keep a non-negligible fraction of samples with no frame references since there might be, in the evaluation benchmarks, questions which do not require CoF-like reasoning, and we do not want to force the model to generate it when unnecessary. This yields a total of 164,186 samples, comprising 103,683 samples from the CoF-Datareal dataset, which is based on real-world videos, and 60,503 samples from the CoF-Datasynth dataset of synthetic videos. Fig. 4 shows the distribution of how many frames are referenced in the reasoning traces, both for the final dataset and the individual splits. The CoF-Datasynth exhibits a more balanced distribution compared to the automatically generated CoF-Datareal: this highlights that using synthetic videos allows us to better control various aspects of the data.
Figure 5: Video encoding format. InternVL models add a textual identifier before each frame, which is well-suited for our chain-of-frames reasoning. The illustration is taken from Chen et al. (2024).
4 Experiments
4.1 Experimental setup
Models selection. While chain-of-frames is a general approach, we find it particularly well-suited for the recent InternVL models (Chen et al., 2024; Zhu et al., 2025). These models introduce a novel format for videos, where frames are interleaved with text identifiers such as Frame-1, Frame-2, etc. (see Fig. 5), reinforcing the temporal structure of the video. This format already associates the image with their corresponding textual identifiers, and using these identifiers in reasoning traces may facilitate long-range interaction in long-context LLMs. The InternVL models achieve state-of-the-art performance among open-weight video LLMs, and thus improving their reasoning capabilities is both challenging and of practical relevance. Consequently, we adopt the recently released InternVL2.5-4B (Chen et al., 2024) and InternVL3-8B (Zhu et al., 2025) models for our experiments. We leave exploration of CoF on other architectures to future work.
Training details. For InternVL2.5-4B, we fully fine-tune both the LLM and projection modules, while keeping the vision encoder frozen. For InternVL3-8B, we apply LoRA-based fine-tuning (Hu et al., 2022) to reduce memory usage. Detailed training configurations can be found in App. A.
Video benchmarks. We compare the video LLMs on five popular benchmarks that capture diverse aspects of video understanding. These benchmarks span a broad range of tasks, video types and duration, providing a comprehensive evaluation of model capabilities. Video-MME (Fu et al., 2024) includes six visual domains, including videos from 2 to 60 minutes long. MVBench (Li et al., 2023b) encompasses 20 tasks that require more than single-frame analysis. VSI-Bench (Yang et al., 2024) focuses on quantitative reasoning tasks such as object counting or appearance order. To evaluate hallucination tendencies of video LLMs, we additionally include the VidHal (Choong et al., 2024) and EventHallusion (Zhang et al., 2024a) benchmarks. Details on each benchmark and the breakdown of the results over the fine-grained splits are available in App. A and App. B.
Table 1: Chain-of-Frames vs other Chain-of-Thoughts variants. We compare different approaches to encourage reasoning in video LLMs, via either prompting or supervised fine-tuning (SFT), see Sec. 4.2 for details. All models are obtained from InternVL2.5-4B. Fine-tuning on our chain-of-frames (CoF) data yields the best accuracy on all benchmarks.
Model VSI-Bench Video-MME MVBench VidHal EventHall
Original 31.8 54.9 70.8 74.0 62.5
Original + CoT Prompting 33.5 54.7 71.5 77.0 67.4
SFT with QA only 31.8 54.5 73.4 64.1 57.7
SFT with CoT 34.3 58.6 73.7 77.9 53.1
SFT with CoF (ours) 36.9 59.7 76.1 79.2 71.2
4.2 Chain-of-Frames vs other Chain-of-Thoughts variants
Setup. We aim to evaluate how our chain-of-frames approach compares to alternative methods for incorporating reasoning into video LLMs, including both prompting and fine-tuning baselines. To ensure consistency, we use InternVL2.5-4B as a baseline model, and for variants requiring fine-tuning, we adopt the training scheme detailed above. We compare the following models.
  • Original: the InternVL2.5-4B model with default prompting.
  • Original + CoT Prompting: the InternVL2.5-4B model with a prompt that encourages the model to perform intermediate reasoning before answering the question (see prompt in App. A). This approach is similar to the standard CoT used in language tasks (Wei et al., 2022).
  • SFT with QA only: the InternVL2.5-4B model fine-tuned on the question-answers pairs from our CoF-Data without including the reasoning traces.
  • SFT with CoT: the InternVL2.5-4B model fine-tuned on CoF-Data, where reasoning traces are included but references to specific frames are removed (e.g., “In Frame 1…” is replaced with a generic “In the video…”). This approach mimics the standard CoT format and is similar to the method of Fei et al. (2024). As their model is not publicly available, this method provides a direct comparison to their approach (VideoCoT) on our more comprehensive dataset.
  • SFT with CoF: the InternVL2.5-4B model fine-tuned on CoF-Data, i.e. our proposed approach.
For all baseline models based on supervised fine-tuning (SFT), we report results using the best prompting strategy (either standard or CoT) for each benchmark. For our SFT with CoF model, we always use CoT prompting across all benchmarks. A complete comparison is provided in App. B.
Results. We report results for all models on the five benchmarks in Table 1. First, we observe that CoT prompting alone already improves the accuracy of the original model on four out of five benchmarks compared to standard prompting. This result highlights the value of encouraging explicit reasoning, which video LLMs are capable of performing. Second, fine-tuning on question-answer pairs without reasoning (SFT with QA only) gives mix results, possibly due to overfitting on the training data, which might degraded the reasoning ability of the original InternVL2.5-4B. Training on reasoning traces without temporal grounding (SFT with CoT) improves the results across nearly all benchmarks, with the notable exception of EventHallusion. Finally, the model trained on the full CoF-Data (SFT with CoF), i.e., fine-tuned on the reasoning traces including frame references, achieves the highest accuracy across all benchmarks, with improvements ranging from 4.8% to 8.7% over the original InternVL2.5-4B. These results demonstrate that our chain-of-frames approach, which combines reasoning with temporal grounding, offers consistent benefits across diverse tasks and domains, from quantitative reasoning to hallucination detection. Finally, as shown in App. C, the reasoning with frame references provide interpretable insights into the model’s decision-making process.
4.3 Comparison to state-of-the-art video LLMs
We want now to contextualize the performance of our CoF-based models with the current leading video LLMs. For this, in Table 2 we report the performance of some of the strongest both closed-source (GPT-4V/4T (Achiam et al., 2023), GPT-4o (Hurst et al., 2024), Gemini-1.5-Pro (Gemini Team et al., 2024)) and open-source (Qwen2-VL-72B (Wang et al., 2024a), LLaVA-OneVision-72B (Li et al., 2024a)) models of different size. We could not add the original models from other works (Wang et al., 2024b; Han et al., 2024; Fei et al., 2024; Hu et al., 2025b) as these are not publicly available. However, we provide comparisons based on the common benchmarks in App. B, where our CoF-models outperform these baselines. Besides the InternVL2.5-4B fine-tuned on CoF-Data, we provide an additional CoF-based LLMs obtained by fine-tuning the more recent InternVL3-8B: we indicate them as CoF-InternVL2.5-4B and CoF-InternVL3-8B respectively. We note that we do not refer to the results of any models on our benchmarks as zero-shot. In fact, for closed-source and some open-source models, the training data is not known. Moreover, 5 out 20 tasks of MVBench are partially based on the test set of Clevrer and VidHal includes some videos from MVBench, and, e.g., InternVL2.5 models have been pre-trained on Clevrer.
Table 2 shows that, despite being significantly smaller than most baselines, CoF-InternVL3-8B achieves the best results on VSI-Bench, Video-MME, and MVBench, outperforming leading (closed-source) video LLMs. Moreover, it has second-best accuracy on VidHal, and best among open-source LLMs on EventHallusion. Meanwhile, CoF-InternVL2.5-4B attains competitive results, surpassing several larger open-source models and in some cases even closed-source ones. These results demonstrate the effectiveness of using chain-of-frames reasoning to improve performance across diverse and challenging video understanding benchmarks. Finally, in App. B we report the breakdown of the results over the subsets of the benchmarks in Table 2. From those we can see, for example, that training with CoF leads to improvements in the spatial reasoning tasks of VSI-Bench. We hypothesize that this is due to having the reasoning traces about object count, appearance order, and relative distance created from Clevrer in CoF-Data. This suggests that the model is effectively learning spatial reasoning tasks from synthetic CoF data.
Table 2: Comparison of CoF-models to state-of-the-art video LLMs. We report accuracy on the five evaluation benchmarks for relevant baseline models and our CoF-models fine-tuned on CoF-Data. Both CoF-InternVL2.5-4B and CoF-InternVL3-8B outperform the majority of baselines despite significantly fewer parameters, and obtain the best performance on three out of five benchmarks.
Model VSI-Bench Video-MME MVBench VidHal EventHall
Closed-source Models
GPT-4V/4T - 59.9 43.7 - 76.5
GPT-4o 34.0 71.9 - 77.2 91.9
Gemini-1.5-Pro 48.8 75.0 - 67.1 80.4
Open-source Models
LLaVA-OneVision-72B 40.2 66.2 59.4 64.7 59.5
Qwen2-VL-72B 37.6 71.2 73.6 76.2 54.7
LLaVA-OneVision-7B 32.4 58.2 56.7 58.4 60.1
LLaVA-NeXT-Video-7B 35.6 46.5 53.1 50.9 64.8
Qwen2-VL-7B 31.0 63.3 67.0 69.6 59.3
InternVL2.5-4B 33.5 54.7 71.5 77.0 67.4
InternVL3-8B 41.0 66.5 74.4 80.9 72.1
Our Models
CoF-InternVL2.5-4B 36.9 59.7 76.1 79.2 71.2
CoF-InternVL3-8B 51.3 75.3 90.4 79.5 78.7
4.4 Additional analyses of Chain-of-Frames
Figure 6: Diversity of training data. Combining CoF examples extracted from real and synthetic videos provides better results on almost all benchmarks compared to using a single source. For all models we use 164k training samples, and fine-tune an InternVL2.5-4B model.
Influence of training data.
The training data of CoF-Data combines chain-of-frames from real and synthetic videos (total of 164k samples). To evaluate the impact of dataset diversity, we construct datasets of the same size (164k samples) but from a single source, i.e., either CoF-Datareal or CoF-Datasynth, and fine-tune InternVL2.5-4B on them. We compare these two models to standard CoF-InternVL2.5-4B (trained on CoF-Data) in Fig. 6. Using the combined dataset outperforms single-source datasets on all benchmarks except for EventHallusion, demonstrating the importance of diversity in the reasoning traces used for training. Between the two individual datasets, the model trained on synthetic videos outperforms the one trained on real videos in three out of five cases, while being worse in just one. This suggests that further improvements might come from expanding the tasks covered by the synthetic data, which can even be generated at low cost.
CoF reasoning at inference time. To better understand how the CoF models utilize frame references during inference, we track the number of frames referenced in the reasoning trace per answer. In Fig. 7, we report how many answers, generated by our CoF-InternVL2.5-4B and CoF-InternVL3-8B models across all evaluation benchmarks, contain $N$ frames for $N=1,\ldots,10$. We can see that both models learn to produce, from a relatively small training set, reasoning traces which refer to frame IDs (see qualitative examples in App. C). In particular, CoF-InternVL2.5-4B includes at least one frame reference in 68.7% of the answers, while CoF-InternVL3-8B does it in 76.9% of the cases. This suggests that larger models might benefit more from fine-tuning with frame-grounded reasoning: in fact, the CoF-InternVL3-8B model shows both higher use of frame references and larger gains in accuracy (see Fig. 1). Finally, the models can use frame references selectively rather than uniformly across tasks ( distribution by benchmark in App. B), depending on the type of questions.
Figure 7: CoF reasoning at inference time. For both CoF-InternVL2.5-4B (left plot) and CoF-InternVL3-8B (right), we show the distribution of the number of frame references generated at inference time: both models learn to produce chain-of-frames during evaluation.
5 Conclusion
We have introduced chain-on-frames (CoF), a new approach to encourage video LLMs to produce temporally grounded reasoning before providing the final answer. Compared to existing works, CoF does not require complex ad-hoc inference frameworks or auxiliary models, and we show that its training data can be efficiently extracted from both real and synthetic videos. Our models fine-tuned on CoF-Data outperform across multiple benchmarks those obtained with alternative methods for reasoning, and even achieve results better than or similar to leading video LLMs. Overall, these features make CoF a viable option to further improve the reasoning capabilities of video LLMs. Exploring the effect of increasing the size and diversity of the training data, as well as the scale of the models, represents an exciting direction for future work.
Limitations. We have currently applied CoF on the InternVL models since their video encoding is well-suited for our reasoning with frame references, and they achieve state-of-the-art performance. It is an interesting open question how to customize our approach to other types of video LLMs.
Acknowledgements
This paper is supported in part by the Army Research Office under grant number W911NF-21-1-0155 and by the New York University Abu Dhabi (NYUAD) Center for Artificial Intelligence and Robotics, funded by Tamkeen under the NYUAD Research Institute Award CG010. Additional support was provided by the NYU IT High Performance Computing resources, services, and staff expertise. Moreover, F.C. and N.F. acknowledge support from an unrestricted gift from Google and by the Swiss National Science Foundation (grant number 212111).
Appendix A Experimental Details
A.1 Chain-of-Frames training data
CoF from real videos (CoF-Datareal). To generate question-reasoning-answer triplets, we prompt Llama-3.1-8B-Instruct (Meta AI, 2024) using the instruction shown in Fig. 8 along with frame-aware video captions from the VideoEspresso dataset (see Fig. 3 for details). Notably, the raw video content is not included in this process. Two examples from CoF-Datareal are shown in Fig. 12.
CoF from synthetic videos (CoF-Datasynth). The second portion of our training dataset is derived from the Clevrer dataset, which includes detailed attributes for each object in every video frame. Specifically, given a frame ID and object ID, the inside_camera_key field indicates whether the object is visible in the frame, enabling us to determine when an object enters or exits the scene. The velocity attribute reflects whether an object is moving or stationary, while the location attribute provides its absolute or relative position, which can be leveraged to estimate distances or identify collisions. The final CoF-Datasynth dataset comprises three categories of questions: Object Count, appearance order, and relative distance. Within the object count category, we define three subtypes: (i) collision-based (“How many collisions…”), (ii) motion state (“How many moving objects…”), and (iii) temporal-based, where questions reference specific segments of the video (“After object A enters…”). The questions, answers and reasoning traces are generated with the manually written templates shown in Fig. 9, making the data collection process particularly simple and fast. Examples from CoF-Datasynth are shown in Fig. 13.
Prompt
Figure 8: Prompt for CoF-Datareal. We prompt Llama-3.1-8B to generate questions, answers, and reasoning traces with reference frames from the real videos of VideoEspresso. Notably, to generate our training data, we do not use the videos but only their captions.
Object Count Template Appearance Order Template Relative Distance Template
Figure 9: Templates for CoF-Datasynth. To generate questions, answers and reasoning traces with reference frames from the annotations of the synthetic videos of Clevrer we rely on fixed, manually written templates. We create three types of questions (object count, appearance order, relative distance) with different templates.
A.2 Video benchmarks
Video-MME. Video-MME (Fu et al., 2024) offers a diverse range of video types, covering six primary visual domains and 30 subfields to support broad scenario generalizability. It also introduces variation in temporal length, including short (under 2 minutes), medium (4-15 minutes), and long (30-60 minutes) videos.
MVBench. Li et al. (2023b) presents a comprehensive benchmark for multimodal video understanding, encompassing 20 challenging tasks that require more than single-frame analysis. It is specifically designed to evaluate a model’s ability to understand temporal dynamics across video sequences.
VSI-Bench. This benchmark (Yang et al., 2024) is designed to quantitatively assess the visual-spatial intelligence of multimodal large language models. Built from over 5,000 high-quality question-answer pairs across 288 real-world indoor videos, VSI-Bench spans diverse environments such as homes, offices, and industrial spaces. The benchmark covers eight tasks: object count, relative distance, relative direction, route planning, object size estimation, room size estimation, absolute distance estimation, and appearance order. Out of these tasks included in this benchmark, relative distance, appearance order, relative directory, and route planning come with multiple-choice questions while the other four require an open-ended quantitative answer. To better evaluate the proximity of the model’s prediction with the correct answer, Yang et al. (2024) proposes using mean relative accuracy ($\mathcal{MRA}$). Given a model’s prediction $\hat{y}$ and ground truth $y$, relative accuracy is calculated by:
$\mathcal{MRA}=\frac{1}{10}\sum_{\theta\in\mathcal{C}}\mathds{1}\left(\frac{|\hat{y}-y|}{y}<1-\theta\right)$
where $\mathcal{C}=\{0.5,0.55,\cdots,0.95\}$ and denotes a range of confidence thresholds $\theta$ to calculate the relative accuracy.
VidHal. To evaluate video-based hallucinations in video LLMs, we use VidHal (Choong et al., 2024), a multiple-choice benchmark that features video instances drawn from public video understanding datasets, covering a diverse array of temporal concepts and aspects such as entity actions and event sequences.
EventHallusion. Zhang et al. (2024a) introduce EventHallusion, from which we use the binary-choice questions designed to systematically assess event-related hallucinations in video LLMs. From a hallucination attribution standpoint, it is specifically curated to evaluate a model’s susceptibility to language priors and vision-language correlation biases.
CoT Prompting
Figure 10: CoT prompt. We show the prompt used for elicit reasoning for both the baseline and our fine-tuned models.
A.3 Chain-of-Frames model
For InternVL2.5-4B, we fully fine-tune both the LLM and the projection modules, keeping the vision encoder frozen. In contrast, for InternVL3-8B, we adopt LoRA-based fine-tuning (Hu et al., 2022) to reduce memory consumption. All other training configurations remain consistent across both models. Training is conducted on a single H100 node equipped with 4 GPUs, using a learning rate of $2\times 10^{-6}$, a batch size of 2, and a single epoch.
Appendix B Additional Experiments
Comparison to additional baselines. We could not include the original models from other works (Wang et al., 2024b; Han et al., 2024; Fei et al., 2024; Hu et al., 2025b) in the main evaluation in Table 2 because they are not accessible to the public, and they do not report results on the five benchmarks we use. However, we provide comparative analyses to M-LLM (Hu et al., 2025b) and Video-of-Thought (Fei et al., 2024) on Video-MME and NextQA (Xiao et al., 2021) using the results they report. In Table 3, we see that our CoF-based models, including CoF-InternVL2.5-4B which has fewer parameters than the competitors, outperform these baselines on both benchmarks. Moreover, compared to the most recent model, M-LLM, our method achieves a greater improvement (4.9% vs 0.8%) on NextQA despite starting from a stronger base model.
Table 3: Comparison to additional baselines. Since the models from Hu et al. (2025b) and Fei et al. (2024) are not publicly available, we compare the results they report on Video-MME and NextQA to those of our CoF models. We also report the video LLM each model is based on and its performance, as well as the improvement of each approach over the original model. Both our CoF models outperform the existing approaches on the two benchmarks.
Backbone Model Video-MME NextQA
Video-LLaVA-7B Original 39.9 66.3
Video-of-Thought (Fei et al., 2024) - 76.0  9.7$\uparrow$
Qwen2-VL-7B Original 58.1 77.6
M-LLM (Hu et al., 2025b) 58.7  0.6$\uparrow$ 78.4  0.8$\uparrow$
InternVL2.5-4B Original 54.9 75.3
SFT with CoF (ours) 59.7  4.8$\uparrow$ 79.6  4.3$\uparrow$
InternVL3-8B Original 66.5 82.4
SFT with CoF (ours) 75.38.8$\uparrow$ 87.34.9$\uparrow$
Effect of CoT prompting. An extended version of Table 1 is presented in Table 4. For all baselines, we report results using two prompting strategies, either standard (indicated by $\bm{\star}$) or chain-of-thought (indicated by $\clubsuit$, the prompt is shown in Fig. 10). For our SFT with CoF models, we always use CoT prompting. When considering InternVL2.5-4B, CoT prompting alone improves the accuracy of the original models on four out of five benchmarks compared to the original model. However, this improvement does not hold for the SFT with QA only variant: we hypothesize that fine-tuning solely on QA data negatively impacts the reasoning capabilities of the baseline model. On the other hand, incorporating reasoning traces into the training data (SFT with CoT) generally enhances the model’s reasoning capabilities, and using CoT prompting is beneficial except for the EventHallusion benchmark. CoT prompting improves the results also for the original InternVL3-8B on most benchmarks. Finally, our models (SFT with CoF) outperform the baseline across all benchmarks.
Table 4: Effect of CoT prompting. For all baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix CoT prompting for our CoF models.
Model Prompt VSI-Bench Video-MME MVBench VidHal EventHall
InternVL2.5-4B
Original $\bm{\star}$ 31.8 54.9 70.8 74.0 62.5
$\clubsuit$ 33.5 54.7 71.5 77.0 67.4
SFT with QA only $\bm{\star}$ 31.8 55.4 70.3 73.6 63.1
$\clubsuit$ 31.8 54.5 73.4 64.1 57.7
SFT with CoT $\bm{\star}$ 31.1 52.6 69.6 74.4 62.5
$\clubsuit$ 34.3 58.6 73.7 77.9 53.1
SFT with CoF (ours) $\clubsuit$ 36.9 59.7 76.1 79.2 71.2
InternVL3-8B
Original $\bm{\star}$ 41.0 62.3 72.0 80.9 72.1
$\clubsuit$ 40.2 66.5 74.3 61.6 73.9
SFT with CoF (ours) $\clubsuit$ 51.3 75.3 90.4 79.5 78.7
Detailed results over benchmark splits. For completeness, we report the fine-grained results over the various splits of VSI-Bench (Table 5), Video-MME (Table 6), MVBench (Table 7), and EventHallusion (Table 8). Moreover, for the baseline models, we report results using two prompting strategies, i.e., standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$).
Detailed results of CoF reasoning at inference time. In Fig. 11, we show statistics of how many frames are referenced in the reasoning traces generated by CoF-InternVL3-8B. To complement Fig. 7, we report the frequency for each benchmark separately. We see that the number of frames mentioned varies across benchmarks, e.g., the cases where no frames are referenced significantly decreases on the hallucination benchmarks VidHal and EventHallusion. This suggests that our CoF models learn to modulate the reasoning traces and the frame references depending on the task.
Figure 11: CoF reasoning at inference time. For each benchmark, we show the frequency of the number of frames referenced in the reasoning traces of CoF-8B.
Table 5: Detailed results on the VSI-Bench benchmark. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
Model Prompt Obj. Count Abs. Dist. Obj. Size Room Size Rel. Dist. App. Order Rel. Dir. Route Plan Avg
InternVL2.5-4B
Original $\bm{\star}$ 29.2 31.2 45.5 20.4 35.4 23.2 41.4 27.8 31.8
$\clubsuit$ 36.0 17.1 38.1 29.8 34.2 30.3 52.2 29.9 33.5
SFT with QA only $\bm{\star}$ 22.7 30.7 44.0 26.0 36.3 22.2 39.4 33.0 31.8
$\clubsuit$ 34.9 18.6 38.8 23.2 37.0 28.2 47.1 26.3 31.8
SFT with CoT $\bm{\star}$ 31.5 22.0 41.6 27.9 36.8 21.2 40.4 27.3 31.1
$\clubsuit$ 39.1 19.5 36.1 26.9 36.1 30.1 57.1 29.4 34.3
SFT with CoF (ours) $\clubsuit$ 42.5 20.8 36.4 29.4 35.4 32.4 62.2 36.1 36.9
InternVL3-8B
Original $\bm{\star}$ 58.6 28.5 49.5 43.3 47.0 38.5 31.3 31.4 41.0
$\clubsuit$ 55.2 29.5 38.1 32.6 42.8 47.1 47.5 28.4 40.2
SFT with CoF (ours) $\clubsuit$ 61.8 34.2 37.7 26.7 66.8 43.4 83.9 55.7 51.3
Table 6: Detailed results on the Video-MME benchmark. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
Model Prompt Short (900) Medium (900) Long (900) Avg
InternVL2.5-4B
Original $\bm{\star}$ 64.9 52.7 47.2 54.9
$\clubsuit$ 64.0 53.2 47.0 54.7
SFT with QA only $\bm{\star}$ 68.0 53.6 44.8 55.5
$\clubsuit$ 66.8 53.1 43.6 54.5
SFT with CoT $\bm{\star}$ 64.3 51.8 41.8 52.6
$\clubsuit$ 70.4 55.7 49.6 58.6
SFT with CoF (ours) $\clubsuit$ 73.1 56.2 49.9 59.7
InternVL3-8B
Original $\bm{\star}$ 73.0 61.7 52.1 62.3
$\clubsuit$ 75.3 65.3 59.0 66.6
SFT with CoF (ours) $\clubsuit$ 80.9 74.4 70.4 75.3
Table 7: Detailed results on the MVBench benchmark. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
Model Prompt AA AC AL AP AS CO CI EN FA MA
InternVL2.5-4B
Original $\bm{\star}$ 89.5 54.0 44.0 75.0 82.9 62.5 79.0 29.0 46.0 97.5
$\clubsuit$ 88.5 50.0 46.5 77.0 81.4 67.0 78.0 34.5 60.5 99.0
SFT with QA only $\bm{\star}$ 90.0 55.0 44.0 76.5 81.9 63.5 75.0 33.0 46.0 98.5
$\clubsuit$ 90.5 50.0 48.5 78.5 84.0 67.5 80.0 38.5 71.5 99.5
SFT with CoT $\bm{\star}$ 87.0 53.0 35.5 76.5 81.4 62.0 76.5 31.5 43.5 98.0
$\clubsuit$ 90.5 46.5 57.5 85.0 83.5 67.0 80.5 38.5 73.5 98.5
SFT with CoF (ours) $\clubsuit$ 93.0 41.0 62.0 91.5 89.4 73.5 79.5 47.0 83.0 98.5
InternVL3-8B
Original $\bm{\star}$ 90.0 42.0 44.5 83.0 82.5 75.5 78.0 38.5 45.0 98.0
$\clubsuit$ 77.5 45.0 42.5 87.5 97.9 80.5 89.0 33.5 54.0 99.0
SFT with CoF (ours) $\clubsuit$ 96.5 95.5 58.0 99.5 100.0 99.0 77.5 79.5 87.5 96.5
Model Prompt MC MD OE OI OS ST SC UA Avg
InternVL2.5-4B
Original $\bm{\star}$ 88.5 73.0 96.5 83.5 39.5 92.0 57.5 85.0 70.8
$\clubsuit$ 86.5 72.5 96.0 81.5 40.5 91.5 58.0 78.0 71.5
SFT with QA only $\bm{\star}$ 87.5 75.0 96.5 82.5 41.0 91.5 59.5 85.5 70.3
$\clubsuit$ 86.5 75.5 96.5 86.5 42.0 92.0 52.5 82.0 73.4
SFT with CoT $\bm{\star}$ 89.0 72.5 96.5 82.0 38.0 91.5 56.5 82.0 69.6
$\clubsuit$ 86.5 72.0 95.5 86.5 40.5 92.0 52.5 80.5 73.7
SFT with CoF (ours) $\clubsuit$ 86.5 72.0 96.5 87.0 44.0 93.5 50.0 82.5 76.1
InternVL3-8B
Original $\bm{\star}$ 60.5 89.0 97.0 85.5 39.5 92.5 69.0 81.0 72.0
$\clubsuit$ 70.5 91.0 97.5 87.5 41.5 92.5 77.0 73.5 74.3
SFT with CoF (ours) $\clubsuit$ 85.0 89.0 90.4 96.5 95.5 94.0 99.0 88.0 90.4
Table 8: Detailed results on the VidHal and EventHallusion benchmarks. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
Model Prompt VidHal EventHallusion
Entire Misleading Mix Avg
InternVL2.5-4B
Original $\bm{\star}$ 74.0 48.3 91.2 48.2 62.5
$\clubsuit$ 77.0 44.7 80.4 77.2 67.4
SFT with QA only $\bm{\star}$ 73.6 48.3 89.2 51.8 63.1
$\clubsuit$ 64.1 47.4 75.5 50.3 57.7
SFT with CoT $\bm{\star}$ 74.4 49.1 91.2 47.1 62.5
$\clubsuit$ 77.9 39.5 71.6 48.2 53.1
SFT with CoF (ours) $\clubsuit$ 79.2 49.1 85.3 79.3 71.2
InternVL3-8B
Original $\bm{\star}$ 80.9 52.6 91.2 72.5 72.1
$\clubsuit$ 61.6 57.0 94.1 70.5 73.9
SFT with CoF (ours) $\clubsuit$ 79.5 57.9 92.2 86.0 78.7
Appendix C Additional Figures
This section presents additional samples from our training dataset along with inference examples. More specifically, Fig. 12 and Fig. 13 show samples from the CoF-Datareal and CoF-Datasynth, respectively. To illustrate the reasoning traces generated by our CoF models and compare them to the answers of the baseline models, we present samples from VSI-Bench and MVBench benchmarks in Fig. 14 and samples from hallucination benchmarks in Fig. 15.
Figure 12: CoF training data. Examples of the training data from CoF-Datareal, including relevant frames, questions, answers, and reasoning traces.
Figure 13: CoF training data. Examples of training data generated from the CoF-Datasynth dataset, including relevant frames, questions, answers, and reasoning traces. The samples shown belong to the object count and relative distance categories, respectively.
(a) VSI-Bench benchmark. We show the question (first box), the answer and possibly CoT reasoning of the original InternVL2.5-4B with CoT prompting (second box), and the answer with CoF reasoning of our CoF-InternVL2.5-4B model (third box).
(b) MVBench benchmark. We show the question (first box), the answer and possibly CoT reasoning of the original InternVL3-8B with CoT prompting (second box), and the answer with CoF reasoning of our CoF-InternVL3-8B model (third box).
Figure 14: Inference examples.
(a) VidHal benchmark. We show the question (first box), the answer, and possibly CoT reasoning of the original InternVL2.5-4B with CoT prompting (second box), and the answer with CoF reasoning of our CoF-InternVL2.5-4B model (third box).
(b) EventHallusion benchmark. We show the question (first box), the answer, and possibly CoT reasoning of the original InternVL3-8B with CoT prompting (second box), and the answer with CoF reasoning of our CoF-InternVL3-8B model (third box).
Figure 15: Inference examples.

### v2

@@
 Recent work has shown that eliciting Large Language Models (LLMs) to generate reasoning traces in natural language before answering the user’s request can significantly improve their performance across tasks.
-This approach has been extended to multimodal LLMs, where the models can produce chain-of-thoughts (CoT) about the content of input images and videos.
-In this work, we propose to obtain video LLMs whose reasoning steps are grounded in, and explicitly refer to, the relevant video frames.
-For this, we first create CoF-Data, a large dataset of diverse questions, answers, and corresponding frame-grounded reasoning traces about both natural and synthetic videos, spanning various topics and tasks.
-Then, we fine-tune existing video LLMs on this chain-of-frames (CoF) data.
-Our approach is simple and self-contained, and, unlike existing approaches for video CoT, does not require auxiliary networks to select or caption relevant frames.
-We show that our models based on CoF are able to generate chain-of-thoughts that accurately refer to the key frames to answer the given question.
-This, in turn, leads to improved performance across multiple video understanding benchmarks, for example, surpassing leading video LLMs on Video-MME, MVBench, and VSI-Bench, and notably reducing the hallucination rate.
-Code available at github.com/SaraGhazanfari/CoF.
+This approach has been extended to multimodal LLMs, where the models can produce chains-of-thoughts (CoT) about the content of input images and videos.
+For video inputs, prior works use complex multi-step pipelines that extract and include relevant frames from videos in the CoT, or produce simpler single-stage reasoning traces at the expense of poor temporal grounding.
+Here, we propose the first video LLMs with single-stage reasoning that includes explicit references to relevant frames, thereby reducing temporal inconsistencies in the reasoning process.
+Our approach is simple, unified, and self-contained, employing a single-stage inference to handle complex video understanding tasks without relying on auxiliary modules for frame selection or caption generation.
+For this, we first create CoF-Data, a large dataset of diverse questions, answers, and corresponding frame-grounded reasoning traces from both natural and synthetic videos, spanning various topics and tasks.
+Our models, obtained fine-tuning video LLMs on this chain-of-frames (CoF) data, generate reasoning traces that accurately identify key frames to answer given questions.
+In turn, this consistently improves performance across multiple video understanding benchmarks.
+Surprisingly, we find that synthetic data alone, despite being out-of-distribution with respect to these real-world benchmarks, provides a significant boost in model accuracy.
+Code available at GitHub.
+(a) A Chain-of-Frames reasoning trace generated by our CoF-4B model.
+(b) CoF models vs the baseline models.
+Figure 1: (a) Chain-of-Frames (CoF) reasoning generated by our CoF-InternVL2.5-4B model, referencing the key frames to answer the question (from Video-MME).
+(b) Comparison of accuracy across multiple video understanding benchmarks between baseline models (InternVL2.5-4B and InternVL3-8B) and their CoF-enhanced counterparts: our models outperform the baselines in nearly every case.
 1 Introduction
-Large Language Models (LLMs) are able to perform step-by-step reasoning, widely known as chain-of-thoughts (CoT) (Wei et al., 2022; Kojima et al., 2022).
-This capability has been implicitly integrated into state-of-the-art systems such as OpenAI’s o1/o3 models (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025), contributing to their remarkable performance and improving the interpretability of their internal functioning.
-CoT reasoning has been also extended to multimodal LLMs (Hu et al., 2024; Wu and Xie, 2023; Shao et al., 2024): this presents new challenges compared to language-only domains, as the models need to attend to inputs from different modalities, and reason about both their individual content and how they are connected (Awal et al., 2023; Kil et al., 2024; Sun et al., 2025).
+Large Language Models (LLMs) are able to perform step-by-step reasoning, widely known as chain-of-thoughts (CoT) [34, 21].
+This capability has been implicitly integrated into state-of-the-art systems such as OpenAI’s o1/o3 models [28] and DeepSeek R1 [13], contributing to their remarkable performance and improving the interpretability of their internal functioning.
+CoT reasoning has been also extended to multimodal LLMs [18, 35, 29]: this presents new challenges compared to language-only domains, as the models need to attend to inputs from different modalities, and reason about both their individual content and how they are connected [3, 20, 30].
 Recent work has also begun to explore the integration of CoT into video understanding tasks, where the input to a multimodal LLM consists of a sequence of images (the frames of the video) along with a text prompt.
 This makes reasoning on videos particularly complex, as the model needs to capture the semantics of the text prompt, understand temporal and causal relationships between frames, and reason about the video in its entirety.
-Existing approaches rely on complex inference frameworks with architecture modifications (Fei et al., 2024) or auxiliary networks (Han et al., 2024) at evaluation time to integrate reasoning into video LLMs.
+Existing approaches rely on complex inference frameworks with architecture modifications [9] or auxiliary networks [14] at evaluation time to extract and integrate key frames from the video sequence into the reasoning trace.
 This makes using these models both more computationally expensive and less general (as they are specialized to some tasks), and deviates from the natural CoT prompting successfully applied to standard LLMs.
@@
 For example, Wang et al.
-(2024b) iteratively refines the samples via annotations from both LLMs and human experts, Han et al.
-(2024) leverages several auxiliary models, Fei et al.
-(2024) uses additional spatial-temporal scene graph data.
+[32] iteratively refine reasoning traces via annotations from LLMs and human experts, Han et al.
+[14] leverage several auxiliary models, and Fei et al.
+[9] use additional spatial-temporal scene graph data.
 Finally, video LLMs do not currently provide explicit connections between particular segments of the video and their reasoning, which would better ground the chain-of-thought.
-To remedy these limitations, in this work we propose chain-of-frames (CoF), a new frame-aware chain-of-thought reasoning approach for video LLMs that integrates temporal information directly into the CoT structure (see Fig.
-1).
-This enables the model to identify and refer to the most relevant frames while answering questions, in contrast to prior works that treat frame selection and reasoning as separate stages.
+To remedy these limitations, in this work we propose Chain-of-Frames (CoF), a new frame-aware chain-of-thought reasoning approach for video LLMs that integrates temporal information directly into a single-shot text-only CoT structure (see Fig.
+1a).
 Chain-of-frames is a simple and natural adaptation of the CoT paradigm in NLP to video understanding that does not require the auxiliary networks or complex inference frameworks of existing methods.
 Moreover, we propose an efficient data generation pipeline that allows us to collect a large dataset of CoF examples, named CoF-Data.
-To achieve this, a key element consists of leveraging a synthetic video dataset (Yi et al., 2020) to extract a large and diverse set of reasoning traces at virtually no cost.
-Then, we fine-tune two recent open-source video LLMs, InternVL2.5-4B (Chen et al., 2024) and InternVL3-8B (Zhu et al., 2025), on our CoF-Data.
-In an extensive evaluation on five established benchmarks, we show that our CoF models significantly outperform the original ones, with and without naive CoT prompting (Wei et al., 2022) (see Fig.
+To achieve this, a key element consists of leveraging a synthetic video dataset [39] to extract a large and diverse set of temporally grounded reasoning traces at virtually no cost.
+Then, we fine-tune three recent open-source video LLMs, InternVL2.5-4B [6], InternVL3-8B [43] and Phi-3.5-Vision [1] on our CoF-Data.
+In an extensive evaluation on five established benchmarks, we show that our CoF models significantly outperform the original ones with and without naive CoT prompting [34] (see Fig.
+1b).
+Moreover, they are competitive with or better than state-of-the-art (SOTA) open- and closed-source models (Tab.
 1).
-Moreover, they are competitive with or better than state-of-the-art (SOTA) open- and closed-source models (Table 2).
-Notably, our CoF-InternVL3-8B achieves higher accuracy than the best model reported on Video-MME, MVBench, and VSI-Bench leaderboards, surpassing even proprietary models like GPT-4o and Gemini 1.5 Pro.
-These results also demonstrate that video LLMs can learn the ability to produce frame-aware reasoning from a limited number of (synthetic) samples and generalize the behavior to unseen tasks.
-Thus, our work provides a simple and inexpensive approach to adapt the CoT paradigm for video LLMs that improves both performance and interpretability.
-(a) A CoF reasoning trace generated by our CoF-4B model.
-(b) CoF models vs the baseline models.
-Figure 1: (a) Chain-of-frames reasoning generated by our CoF-InternVL2.5-4B model: it includes the key frames to answer the question (from Video-MME).
-(b) Comparison of accuracy across multiple video understanding benchmarks between baseline models (InternVL2.5-4B and InternVL3-8B) and their CoF-enhanced counterparts: our models consistently outperform the baselines.
+Finally, further analysis reveals that synthetic data alone—despite being out-of-distribution relative to real-world benchmarks—provides a substantial improvement in model accuracy.
+This indicates that the model effectively learns the underlying task from the synthetic data and successfully transfers this knowledge to real-world scenarios.
 Figure 2: CoF-Data.
-We show examples of our training data with chain-of-frames reasoning, including video, question, answer and reasoning trace.
+We show examples of our training data with chain-of-frames reasoning, including video, question, reasoning trace and answer.
 We include samples from CoF-Datareal (real video, top row) and CoF-Datasynth (synthetic video, bottom), created as described in Sec.
-3.3.
+$3.$
 2 Related Work
 Multimodal LLMs for videos.
-Multimodal Large Language Models (Chen et al., 2023b; Li et al., 2024b; Ghazanfari et al., 2024) have made substantial progress in integrating visual and textual modalities, enabling them to perform complex reasoning and achieve deep understanding across various data types, including videos (Chen et al., 2023b; Xue et al., 2024; Li et al., 2023a; Zhang et al., 2024b, c; Li et al., 2024a; Bai et al., 2024; Zhu et al., 2025).
-Among recent advancements, InternVL2.5 (Chen et al., 2024), InternVL3 (Zhu et al., 2025), LLaVA-NeXT-Video (Zhang et al., 2024c), and Qwen2-VL (Wang et al., 2024a) stand out for their strong video understanding capabilities.
+Multimodal Large Language Models [7, 25, 12] have made substantial progress in integrating visual and textual modalities, enabling them to perform complex reasoning and achieve deep understanding across various data types, including videos [7, 37, 22, 41, 42, 23, 4, 43].
+Among recent advancements, InternVL2.5 [6], InternVL3 [43], LLaVA-NeXT-Video [42], and Qwen2-VL [31] stand out for their strong video understanding capabilities.
 Both InternVL and LLaVA-NeXT-Video process videos in an image-text interleaved format, aligning sequences of video frames with language to form a unified multimodal stream.
-In addition to these open-source models, closed-source systems such as GPT-4o (Achiam et al., 2023) and Gemini-1.5 (Gemini Team et al., 2024) have also demonstrated impressive multimodal capabilities, although details of their architecture and training remain proprietary.
+In addition to these open-source models, closed-source systems such as GPT-4o [2] and Gemini-1.5 [11] have also demonstrated impressive multimodal capabilities, although details of their architecture and training remain proprietary.
 Chain-of-Thoughts for videos.
-Recent research on chain-of-thoughts for video understanding broadly falls into two categories: methods that provide explicit textual reasoning explanations (Wang et al., 2024b), and those that focus on identifying relevant frames to facilitate the process of generating the response (Han et al., 2024; Hu et al., 2025b, a).
-In the first category, VideoCoT (Wang et al., 2024b) introduces an active annotation tool to generate reasoning explanations, thereby encouraging models to explicitly reason on visual content.
-In the second category, several approaches emphasize frame selection as a pre-processing step to enhance reasoning efficiency.
-M-LLM (Hu et al., 2025b) proposes leveraging multimodal LLMs to identify the most relevant frames corresponding to the query.
-Likewise, Chain-of-Shot (Hu et al., 2025a) introduces a prompting strategy specifically tailored for understanding long-form videos by selecting key frames.
-VideoEspresso (Han et al., 2024) combines the ideas of core frame selection with fine-grained reasoning annotations, creating a large-scale dataset that supports more efficient and focused video reasoning.
-Finally, Fei et al.
-(2024) introduces Video-of-Thought, a complex five-step pipeline to generate spatial-temporal scene graphs to answer multiple-choice questions about videos.
+Recent research on chain-of-thought (CoT) reasoning for video understanding can be broadly categorized into two paradigms: (1) single-stage methods, which directly generate textual reasoning explanations followed by an answer in a single step [32], and (2) multi-stage methods, which employ auxiliary models to extract meta-information in earlier stages and then use this information to generate the answer in the final step [14, 17, 16].
+In the first category, VideoCoT [32] introduces an active annotation framework to produce reasoning explanations from video descriptions, thereby guiding the model to generate reasoning traces followed by the answer.
+In the second category, most approaches [33, 17, 16, 14, 5] emphasize key frame selection as an initial stage to enhance downstream answer generation.
+Building on this idea, VideoEspresso [14] further integrates additional steps, combining the selected core frames with the question and feeding them into a reasoning model that performs chain-of-thought (CoT) reasoning, with the final answer produced in a subsequent stage.
+Finally, Video-of-Thought [9] introduces a more complex five-stage pipeline that constructs spatio-temporal scene graphs to reason over videos and answer multiple-choice questions.
 3 Chain-of-Frames: Reasoning on Videos via Frame References
 3.1 Limitations of reasoning on videos
-Multimodal LLMs (Zhang et al., 2024c; Chen et al., 2024; Bai et al., 2024) process videos as a sequence of images (frames), which are encoded as image tokens by a vision encoder, then concatenated with the user prompt in natural language, and finally passed to a language model.
+Recent work has begun to explore the integration of CoT into video understanding tasks.
 While the base language models may be trained to produce reasoning traces, these are not specific for reasoning on videos.
 To encourage chain-of-thought output in video LLMs, models must be fine-tuned on video-grounded reasoning traces.
-Wang et al.
-(2024b) introduces VideoCoT, which used reasoning traces generated by LLMs and refined by human experts to describe video events.
-However, the traces lack explicit temporal grounding, that is, individual reasoning steps are not clearly aligned with the corresponding video frames.
-Additionally, the reliance on human and LLM annotations makes data generation expensive, limiting the dataset to only 11k samples.
-A different strategy is proposed by Han et al.
-(2024), who use multiple auxiliary models to generate CoT data.
-Lightweight multimodal LLMs select the core frames, GPT-4o (Hurst et al., 2024) identifies key elements in each core frame and their relevance to the query, GroundingDINO (Liu et al., 2024) provides spatial annotations (e.g., bounding boxes), and the BGE-M3 retriever (Chen et al., 2023a) produces the temporal annotations.
-While effective during training, this approach cannot be deployed at inference time given its high complexity and cost (only the frame selection step, which depends on auxiliary LLMs, is kept).
-Alternatively, Video-of-Thought (Fei et al., 2024) introduces a complex five-step inference pipeline, which includes generating the spatial-temporal scene graphs for key frames and is specialized for multiple-choice questions.
-In summary, existing approaches to video reasoning with LLMs face three main limitations: (i) expensive training data generation, (ii) complex inference, possibly involving auxiliary models, and (iii) lack of explicit temporal grounding in the reasoning process.
+VideoCoT [32] addresses this by generating CoT training data through a pipeline that leverages video descriptions as raw annotations and uses LLMs to produce corresponding reasoning traces.
+However, because the descriptions are not temporally aligned at the frame level, the resulting reasoning traces can introduce temporal inconsistencies [24].
+To mitigate such inaccuracies, human annotators are introduced into the loop to refine the LLM-generated reasoning traces.
+As a result, the reliance on human and LLM annotations makes data generation expensive, limiting the dataset to only 11k samples.
+Moreover, the traces lack explicit temporal grounding; that is, individual reasoning steps are not clearly aligned with the corresponding video frames.
+A different strategy is to leverage multiple auxiliary models to generate the meta-information, including finding the key frames from the video and explicitly grounding the model’s answer to the key frames [14, 17, 9].
+While these multi-stage pipelines improve video reasoning, their computational overhead during inference poses a major bottleneck.
+Furthermore, because only a subset of the frames is passed to the video LLM, the video’s full temporal context is lost.
+In summary, current video reasoning methods face two major limitations: (1) they depend on expensive and often inaccurate training data generation while lacking explicit temporal grounding in the reasoning process, or (2) involve complex inference procedures that rely on auxiliary models and compromise contextual understanding by processing only a subset of the frames to generate the final answer.
+Figure 3: Overview of our two-step pipeline for generating CoF-Data.
+Step 1 adjusts the frame IDs while preserving frame-caption alignment.
+Step 2 utilizes raw annotations to generate CoF triplets (question, frame-aware reasoning trace, answer), using Llama-3.1-8B-Instruct for real-video data and manual templates for synthetic data.
 3.2 Chain-of-Frames
 To address these limitations, we propose Chain-of-Frames (CoF), a simple yet effective approach introducing temporal grounding into the reasoning process.
-CoF consists of reasoning traces with explicit references to frames relevant to answering the given query.
+CoF consists of reasoning traces with explicit references to frames relevant to answering the given query, within a single-stage inference.
 Concretely, we use the position of the frame in the video (e.g., “Frame 1”, “Frame 2”, …) as an identifier.
@@
 C.
-This approach has several benefits
+This approach has several benefits:
   1. 1.
-    Data availability: as we show in Sec.
-3.3, training data can be efficiently generated from existing annotated real videos.
-Additionally, CoF traces can be created from synthetic videos at virtually no cost.
-This low-cost data generation contrasts with the more complex pipelines of Wang et al.
-(2024b); Han et al.
-(2024).
+    Data quality: as we show in Sec.
+3.3, training data can be efficiently generated from existing annotated real videos, where the video descriptions are temporally aligned at the frame level, resulting in temporally accurate reasoning traces.
+Moreover, CoF traces can be generated from synthetic videos with high precision by combining their rich annotations with manually designed templates, at virtually no additional cost.
+This low-cost and scalable data generation process stands in contrast to the more complex pipelines [32, 14].
   2. 2.
-    Simplicity: the reasoning traces are completely in natural language (without the need for complex format, bounding boxes or scene graphs as for Fei et al.
-(2024); Han et al.
-(2024)).
-This approach is a natural extension of standard CoT method used in NLP tasks and does not require specific inference pipelines.
+    Simplicity: our reasoning traces are expressed entirely in natural language—without requiring auxiliary modules to generate complex data formats, bounding boxes, or scene graphs as in Fei et al.
+[9], Han et al.
+[14].
+This makes our approach a natural extension of standard Chain-of-Thought (CoT) reasoning used in NLP tasks, eliminating the need for multi-stage inference pipelines and reducing overall latency.
   3. 3.
-    Temporal grounding: by explicitly referencing the frames relevant to the task, CoF increases the connection between video and reasoning.
-Thus, our method can be seen as a temporally grounded extension of the typical CoT of Wang et al.
-(2024b).
+    Explicit temporal grounding: temporal grounding is achieved by explicitly referencing the task-relevant frames within the reasoning traces, rather than by retrieving and processing key frames through a multi-stage inference process.
+This design strengthens the connection between visual content and reasoning, making CoF a temporally grounded extension of the typical CoT approach proposed by Wang et al.
+[32].
   4. 4.
-    Interpretability: the reasoning traces generated at inference time include frame references (Fig. 1) and provide direct insights into how the LLM’s decisions are obtained.
-In the next section, we detail how we generate chain-of-frames from video datasets. These traces constitute the training data for fine-tuning video LLMs capable of producing CoF-based reasoning.
-Figure 3: Overview of our two-step pipeline for generating CoF-Data.
-Step 1 adjusts the frame IDs while preserving frame-caption alignment.
-Step 2 utilizes raw annotations to generate CoF triplets (question, frame-aware reasoning trace, answer).
-For this, we leverage Llama3.1-8B a manual template with the synthetic videos from Clevrer.
-3.3 Chain-of-Frames training data collection (CoF-Data)
+    Interpretability: The reasoning traces generated at inference time include frame references (Fig. 1), providing direct insights into how the LLM’s decisions are obtained.
+Figure 4: Video encoding format. InternVL models add a textual identifier before each frame, which is well-suited for our chain-of-frames reasoning. The illustration is taken from Chen et al. [6].
+3.3 Chain-of-Frames training data collection
 We construct chain-of-frames traces from both real and synthetic videos.
-For real videos, we use the training split of the VideoEspresso dataset (Han et al., 2024), which features videos from a wide range of sources and includes descriptions of key frames for each.
+For real videos, we use the training split of the VideoEspresso dataset [14], which features videos from a wide range of sources and includes descriptions of key frames for each.
 These diverse frame-level annotations provide a starting point for extracting CoF traces.
-To complement this, we use the training split of the Clevrer dataset (Yi et al., 2020), which contains synthetic videos of simple 3D objects interacting within a controlled environment, along with rich annotations.
-Synthetic data offers two key advantages: it overcomes the cost of generating training samples from real videos and mitigates the scarcity of sufficiently annotated real videos.
-We next describe the main steps of our data generation pipeline (illustrated in Fig.
-3), and introduce the resulting dataset of chain-of-frames, named CoF-Data.
+To complement this, we use the training split of the Clevrer dataset [39], which contains synthetic videos of simple 3D objects interacting within a controlled environment, along with rich annotations.
+Synthetic data offers two key advantages: it overcomes the cost of generating training samples from real videos and provides highly accurate temporally grounded reasoning trace derived from the rich annotations with manual templates.
+We next describe the main steps of our data generation pipeline, illustrated in Fig.
+3.
 Additional details and examples are provided in App.
@@
 3 (raw annotations).
-From these annotations, we generate triplets of questions, answers, and reasoning traces with frame references by prompting an LLM using the raw annotations as input.
-For this, we use Llama3.1-8B-Instruct (Meta AI, 2024) (the full prompt is provided in App.
+From these temporally aligned annotations, we generate triplets of questions, answers, and reasoning traces with frame references by prompting an LLM using the raw annotations as input.
+For this, we use Llama3.1-8B-Instruct [27] (the full prompt is provided in App.
 A).
-This process yields multiple questions per video, often covering diverse parts of the video and referencing different sets of frames.
-An example from CoF-Datareal is illustrated in Fig.
+This process yields multiple questions per video, often covering diverse parts of the video and referencing different subsets of frames.
+An example from CoF-Datareal is in Fig.
 3, and a complete training sample including the video is in Fig.
@@
 An example of an appearance order question is shown in Fig.
-2 (examples from other categories are shown in App.
+2 (for other categories see App.
 C).
-Figure 4: Distribution of frame references in the Chain-of-Frames training data.
-The left pie chart illustrates the distribution for CoF-Datareal, having fewer frames per reasoning trace, whereas CoF-DATA-synth demonstrates a more balanced frame distribution due to controlled synthetic video generation.
-The right pie chart shows the overall distribution for the CoF-Data.
 Final dataset (CoF-Data).
@@
 We nevertheless keep a non-negligible fraction of samples with no frame references since there might be, in the evaluation benchmarks, questions which do not require CoF-like reasoning, and we do not want to force the model to generate it when unnecessary.
-This yields a total of 164,186 samples, comprising 103,683 samples from the CoF-Datareal dataset, which is based on real-world videos, and 60,503 samples from the CoF-Datasynth dataset of synthetic videos.
+The resulting dataset comprises 164,186 samples.
 Fig.
-4 shows the distribution of how many frames are referenced in the reasoning traces, both for the final dataset and the individual splits.
-The CoF-Datasynth exhibits a more balanced distribution compared to the automatically generated CoF-Datareal: this highlights that using synthetic videos allows us to better control various aspects of the data.
-Figure 5: Video encoding format. InternVL models add a textual identifier before each frame, which is well-suited for our chain-of-frames reasoning. The illustration is taken from Chen et al. (2024).
+5 illustrates the distribution of the number of frames referenced per sample.
+Figure 5: Distribution of frame references in CoF-Data.
 4 Experiments
 4.1 Experimental setup
-Models selection.
-While chain-of-frames is a general approach, we find it particularly well-suited for the recent InternVL models (Chen et al., 2024; Zhu et al., 2025).
+Model VSI-Bench Video-MME MVBench VidHal EventHall Average
+Closed-source Models
+GPT-4V/4T - 59.9 43.7 - 76.5 -
+GPT-4o 34.0 71.9 - 77.2 91.9 -
+Gemini-1.5-Pro 48.8 75.0 - 67.1 80.4 -
+Open-source Models
+LLaVA-OneVision-72B 40.2 66.2 59.4 64.7 59.5 58.0
+Qwen2-VL-72B 37.6 71.2 73.6 76.2 54.7 62.7
+LLaVA-OneVision-7B 32.4 58.2 56.7 58.4 60.1 53.2
+LLaVA-NeXT-Video-7B 35.6 46.5 53.1 50.9 64.8 50.2
+Qwen2-VL-7B 31.0 63.3 67.0 69.6 59.3 58.0
+InternVL2.5-4B 33.5 54.7 71.5 77.0 67.4 60.8
+InternVL3-8B 41.0 66.5 74.4 80.9 72.1 67.0
+Our Models
+CoF-InternVL2.5-4B 36.9 59.7 76.1 79.2 71.2 64.6
+CoF-InternVL3-8B 51.3 73.7 77.1 79.5 78.7 72.1
+Table 1: Comparison of CoF-models to state-of-the-art video LLMs.
+We report accuracy on five benchmarks (and their average) for relevant baselines and our CoF-models fine-tuned on CoF-Data.
+Both CoF-InternVL2.5-4B and CoF-InternVL3-8B outperform the majority of baselines despite significantly fewer parameters, and obtain the best or second-best results on four out of five benchmarks.
+Baseline models.
+While chain-of-frames is a general approach, we find it particularly well-suited for the recent InternVL models [6, 43].
 These models introduce a novel format for videos, where frames are interleaved with text identifiers such as Frame-1, Frame-2, etc.
 (see Fig.
-5), reinforcing the temporal structure of the video.
+4), reinforcing the temporal structure of the video.
 This format already associates the image with their corresponding textual identifiers, and using these identifiers in reasoning traces may facilitate long-range interaction in long-context LLMs.
 The InternVL models achieve state-of-the-art performance among open-weight video LLMs, and thus improving their reasoning capabilities is both challenging and of practical relevance.
-Consequently, we adopt the recently released InternVL2.5-4B (Chen et al., 2024) and InternVL3-8B (Zhu et al., 2025) models for our experiments.
-We leave exploration of CoF on other architectures to future work.
+Consequently, we adopt the recently released InternVL2.5-4B [6] and InternVL3-8B [43] models for our experiments.
+To test the generalizability of our approach beyond the InternVL family, we use we employ Phi-3.5-Vision [1], a mobile-scale multimodal LLM that demonstrates strong performance in language reasoning, as a third baseline model.
+For limited spaces, we present the results relative to Phi-3.5-Vision in App.
+B, where we show that CoF is also effective on this video LLM.
 Training details.
 For InternVL2.5-4B, we fully fine-tune both the LLM and projection modules, while keeping the vision encoder frozen.
-For InternVL3-8B, we apply LoRA-based fine-tuning (Hu et al., 2022) to reduce memory usage.
+For InternVL3-8B, we apply LoRA-based fine-tuning [15] to reduce memory usage.
+We refer to the resulting fine-tuned models as CoF-InternVL2.5-4B and CoF-InternVL3-8B throughout the paper.
 Detailed training configurations can be found in App.
 A.
+Inference details.
+We uniformly sample 30 frames from each input video for inference.
+This approach ensures consistent temporal coverage regardless of video length.
+Importantly, the model is not restricted to 30-second videos and can process videos of arbitrary duration.
 Video benchmarks.
-We compare the video LLMs on five popular benchmarks that capture diverse aspects of video understanding.
-These benchmarks span a broad range of tasks, video types and duration, providing a comprehensive evaluation of model capabilities.
-Video-MME (Fu et al., 2024) includes six visual domains, including videos from 2 to 60 minutes long.
-MVBench (Li et al., 2023b) encompasses 20 tasks that require more than single-frame analysis.
-VSI-Bench (Yang et al., 2024) focuses on quantitative reasoning tasks such as object counting or appearance order.
-To evaluate hallucination tendencies of video LLMs, we additionally include the VidHal (Choong et al., 2024) and EventHallusion (Zhang et al., 2024a) benchmarks.
+We compare the video LLMs on five popular benchmarks that span a broad range of tasks, video types and duration, providing a comprehensive evaluation of model capabilities.
+Video-MME [10] includes six visual domains, including videos from 2 to 60 minutes long.
+MVBench [26] encompasses 20 tasks that require more than single-frame analysis.
+VSI-Bench [38] focuses on quantitative reasoning tasks like object counting or appearance order.
+To evaluate hallucination in video LLMs, we include VidHal [8] and EventHallusion [40].
 Details on each benchmark and the breakdown of the results over the fine-grained splits are available in App.
@@
 B.
+4.2 Main experiments
+We now contextualize the performance of our CoF-based models in relation to the current leading video LLMs.
+For this, in Tab.
+1 we report the performance of some of the strongest both closed-source (GPT-4V/4T [2], GPT-4o [19], Gemini-1.5-Pro [11]) and open-source (Qwen2-VL-72B [31], LLaVA-OneVision-72B [23]) models of different sizes.
+We could not add the prior video CoT models [32, 14, 9, 17] in Tab.
+1 as these are not publicly available.
+Nevertheless, Tab.
+3 shows comparisons on shared benchmark datasets, where our CoF-based models outperform these baselines.
+Finally, we note that we do not refer to the results of any models on our benchmarks as zero-shot.
+In fact, for closed-source and some open-source models, the training data is not known.
+Moreover, 5 out 20 tasks of MVBench are partially based on the test set of Clevrer and VidHal includes some videos from MVBench.
+Comparison to base models.
+Tab.
+1 shows that CoF-enhanced models consistently outperform their corresponding base models across nearly all benchmarks.
+CoF-InternVL3-8B achieves 5.1% higher average performance than the original InternVL3-8B, while the CoF-InternVL2.5-4B improves of 3.8% over InternVL2.5-4B.
+Moreover, the fact that performance gains are more notable for the larger model suggests that more capable LLMs may benefit more substantially from CoF-based fine-tuning.
+Comparison to state-of-the-art video LLMs.
+Tab.
+1 also illustrates that, despite being significantly smaller than most baselines, CoF-InternVL3-8B achieves the best results on VSI-Bench and MVBench, outperforming leading (closed-source) video LLMs.
+It also has second-best accuracy on Video-MME and VidHal, and best among open-source LLMs on EventHallusion.
+Meanwhile, CoF-InternVL2.5-4B attains competitive results, surpassing several larger open-source models and in some cases even closed-source ones.
+These results demonstrate the effectiveness of chain-of-frames reasoning in improving video understanding across diverse and challenging benchmarks.
+Comparison to video CoT methods.
+Prior video CoT models [32, 14, 9, 17] are not accessible to the public, and they do not report results on the five benchmarks we use.
+However, we provide comparative analyses to M-LLM [17] and Video-of-Thought [9] on Video-MME and NextQA [36] using the results they report.
+Both baselines adopt multi-stage inference that explicitly retrieves key frames, whereas our approach integrates frame retrieval implicitly through frame references within the reasoning traces.
+In Tab.
+3, we see that our CoF-based models, including CoF-InternVL2.5-4B, which has fewer parameters than the competitors, outperform these baselines on both benchmarks.
+Moreover, compared to the most recent model, M-LLM, our method achieves a greater improvement (4.9% vs 0.8%) on NextQA despite starting from a stronger base model.
+5 Additional analyses
+5.1 Influence of synthetic training data
+Figure 6: Influence of synthetic training data.
+Across most benchmarks, the model trained on synthetic data consistently outperforms the one trained on real data.
+Despite the substantial distribution gap between the synthetic samples and the real frames in these benchmarks, the model effectively learns the tasks from synthetic data and transfers this knowledge to real-world scenarios.
+For all models we use 164k training samples, and fine-tune an InternVL2.5-4B model.
+The training data of CoF-Data combines chain-of-frames from real and synthetic videos.
+As mentioned in Sec.
+3.3, the synthetic portion of the dataset, CoF-Datasynth, is generated without relying on heavy models, but rather through human-designed templates.
+The key question is how effective this synthetic data is for model training.
+To address this question, we construct datasets of equal size, each sourced exclusively from either CoF-Datareal or CoF-Datasynth, and fine-tune InternVL2.5-4B on them.
+We then compare the resulting CoF-InternVL2.5-4B models in Fig.
+6.
+First, we observe that across most benchmarks, the model trained on synthetic data consistently outperforms the one trained on real data.
+This result is notable: despite the considerable distributional gap between the synthetic samples and real frames from the benchmarks, the model effectively learns tasks such as object counting and appearance ordering from synthetic examples.
+Since synthetic data can be generated at low cost and large scale, this finding highlights a promising and scalable strategy for improving performance.
+Second, using the combined dataset outperforms single-source datasets on all benchmarks except for EventHallusion, demonstrating the importance of diversity in the reasoning traces used for training.
+5.2 Chain-of-frames vs chain-of-thoughts variants
+Backbone Model Video-MME NextQA
+Video-LLaVA-7B Original 39.9 66.3
+[…]
+Qwen2-VL-7B Original 58.1 77.6
+[…]
+InternVL2.5-4B Original 54.9 75.3
+[…]
+InternVL3-8B Original 66.5 82.4
+[…]
+Table 2: Comparison to video CoT methods.
+Since the models from Hu et al.
+[17] and Fei et al.
+[9] are not publicly available, we compare the results they report on Video-MME and NextQA to those of our CoF models.
+We also report the improvement of each approach over their baseline.
+Both our CoF models outperform the existing approaches on the two benchmarks.
+Model VSI-Bench Video-MME MVBench VidHal EventHall
+Original 31.8 54.9 70.8 74.0 62.5
++ CoT Prompting 33.5 54.7 71.5 77.0 67.4
++ SFT with QA only 31.8 54.5 73.4 64.1 57.7
++ SFT with CoT 34.3 58.6 73.7 77.9 53.1
++ SFT with CoF (ours) 36.9 59.7 76.1 79.2 71.2
 Table 3: Chain-of-Frames vs other Chain-of-Thoughts variants.
@@
 Fine-tuning on our chain-of-frames (CoF) data yields the best accuracy on all benchmarks.
-Model VSI-Bench Video-MME MVBench VidHal EventHall
-Original 31.8 54.9 70.8 74.0 62.5
-Original + CoT Prompting 33.5 54.7 71.5 77.0 67.4
-SFT with QA only 31.8 54.5 73.4 64.1 57.7
-SFT with CoT 34.3 58.6 73.7 77.9 53.1
-SFT with CoF (ours) 36.9 59.7 76.1 79.2 71.2
-4.2 Chain-of-Frames vs other Chain-of-Thoughts variants
 Setup.
@@
 We compare the following models.
-  • Original: the InternVL2.5-4B model with default prompting.
+  • Original: InternVL2.5-4B with default prompting.
   • Original + CoT Prompting: the InternVL2.5-4B model with a prompt that encourages the model to perform intermediate reasoning before answering the question (see prompt in App.
 A).
-This approach is similar to the standard CoT used in language tasks (Wei et al., 2022).
-  • SFT with QA only: the InternVL2.5-4B model fine-tuned on the question-answers pairs from our CoF-Data without including the reasoning traces.
-  • SFT with CoT: the InternVL2.5-4B model fine-tuned on CoF-Data, where reasoning traces are included but references to specific frames are removed (e.g., “In Frame 1…” is replaced with a generic “In the video…”).
-This approach mimics the standard CoT format and is similar to the method of Fei et al.
-(2024).
-As their model is not publicly available, this method provides a direct comparison to their approach (VideoCoT) on our more comprehensive dataset.
-  • SFT with CoF: the InternVL2.5-4B model fine-tuned on CoF-Data, i.e. our proposed approach.
+This approach is similar to the standard CoT used in language tasks [34].
+  • Original + SFT with QA only: the InternVL2.5-4B model fine-tuned on the question-answers pairs from our CoF-Data without including the reasoning traces.
+  • Original + SFT with CoT: the InternVL2.5-4B model fine-tuned on CoF-Data, where reasoning traces are included but references to specific frames are removed (e.g., “In Frame 1…” is replaced with a generic “In the video…”).
+  • Original + SFT with CoF: the InternVL2.5-4B model fine-tuned on CoF-Data, i.e. our proposed approach.
 For all baseline models based on supervised fine-tuning (SFT), we report results using the best prompting strategy (either standard or CoT) for each benchmark.
@@
 Results.
-We report results for all models on the five benchmarks in Table 1.
-First, we observe that CoT prompting alone already improves the accuracy of the original model on four out of five benchmarks compared to standard prompting.
+We report results for all models on the five benchmarks in Tab.
+3.
+First, we observe that CoT prompting alone already improves the accuracy of the original model compared to standard prompting.
 This result highlights the value of encouraging explicit reasoning, which video LLMs are capable of performing.
-Second, fine-tuning on question-answer pairs without reasoning (SFT with QA only) gives mix results, possibly due to overfitting on the training data, which might degraded the reasoning ability of the original InternVL2.5-4B.
+Second, fine-tuning on question-answer pairs without reasoning (SFT with QA only) gives mix results, possibly due to overfitting on the training data, which might degraded the reasoning ability of the original model.
 Training on reasoning traces without temporal grounding (SFT with CoT) improves the results across nearly all benchmarks, with the notable exception of EventHallusion.
-Finally, the model trained on the full CoF-Data (SFT with CoF), i.e., fine-tuned on the reasoning traces including frame references, achieves the highest accuracy across all benchmarks, with improvements ranging from 4.8% to 8.7% over the original InternVL2.5-4B.
-These results demonstrate that our chain-of-frames approach, which combines reasoning with temporal grounding, offers consistent benefits across diverse tasks and domains, from quantitative reasoning to hallucination detection.
-Finally, as shown in App.
-C, the reasoning with frame references provide interpretable insights into the model’s decision-making process.
-4.3 Comparison to state-of-the-art video LLMs
-We want now to contextualize the performance of our CoF-based models with the current leading video LLMs.
-For this, in Table 2 we report the performance of some of the strongest both closed-source (GPT-4V/4T (Achiam et al., 2023), GPT-4o (Hurst et al., 2024), Gemini-1.5-Pro (Gemini Team et al., 2024)) and open-source (Qwen2-VL-72B (Wang et al., 2024a), LLaVA-OneVision-72B (Li et al., 2024a)) models of different size.
-We could not add the original models from other works (Wang et al., 2024b; Han et al., 2024; Fei et al., 2024; Hu et al., 2025b) as these are not publicly available.
-However, we provide comparisons based on the common benchmarks in App.
-B, where our CoF-models outperform these baselines.
-Besides the InternVL2.5-4B fine-tuned on CoF-Data, we provide an additional CoF-based LLMs obtained by fine-tuning the more recent InternVL3-8B: we indicate them as CoF-InternVL2.5-4B and CoF-InternVL3-8B respectively.
-We note that we do not refer to the results of any models on our benchmarks as zero-shot.
-In fact, for closed-source and some open-source models, the training data is not known.
-Moreover, 5 out 20 tasks of MVBench are partially based on the test set of Clevrer and VidHal includes some videos from MVBench, and, e.g., InternVL2.5 models have been pre-trained on Clevrer.
-Table 2 shows that, despite being significantly smaller than most baselines, CoF-InternVL3-8B achieves the best results on VSI-Bench, Video-MME, and MVBench, outperforming leading (closed-source) video LLMs.
-Moreover, it has second-best accuracy on VidHal, and best among open-source LLMs on EventHallusion.
-Meanwhile, CoF-InternVL2.5-4B attains competitive results, surpassing several larger open-source models and in some cases even closed-source ones.
-These results demonstrate the effectiveness of using chain-of-frames reasoning to improve performance across diverse and challenging video understanding benchmarks.
-Finally, in App.
-B we report the breakdown of the results over the subsets of the benchmarks in Table 2.
-From those we can see, for example, that training with CoF leads to improvements in the spatial reasoning tasks of VSI-Bench.
-We hypothesize that this is due to having the reasoning traces about object count, appearance order, and relative distance created from Clevrer in CoF-Data.
-This suggests that the model is effectively learning spatial reasoning tasks from synthetic CoF data.
-Table 2: Comparison of CoF-models to state-of-the-art video LLMs.
-We report accuracy on the five evaluation benchmarks for relevant baseline models and our CoF-models fine-tuned on CoF-Data.
-Both CoF-InternVL2.5-4B and CoF-InternVL3-8B outperform the majority of baselines despite significantly fewer parameters, and obtain the best performance on three out of five benchmarks.
-Model VSI-Bench Video-MME MVBench VidHal EventHall
-Closed-source Models
-GPT-4V/4T - 59.9 43.7 - 76.5
-GPT-4o 34.0 71.9 - 77.2 91.9
-Gemini-1.5-Pro 48.8 75.0 - 67.1 80.4
-Open-source Models
-LLaVA-OneVision-72B 40.2 66.2 59.4 64.7 59.5
-Qwen2-VL-72B 37.6 71.2 73.6 76.2 54.7
-LLaVA-OneVision-7B 32.4 58.2 56.7 58.4 60.1
-LLaVA-NeXT-Video-7B 35.6 46.5 53.1 50.9 64.8
-Qwen2-VL-7B 31.0 63.3 67.0 69.6 59.3
-InternVL2.5-4B 33.5 54.7 71.5 77.0 67.4
-InternVL3-8B 41.0 66.5 74.4 80.9 72.1
-Our Models
-CoF-InternVL2.5-4B 36.9 59.7 76.1 79.2 71.2
-CoF-InternVL3-8B 51.3 75.3 90.4 79.5 78.7
-4.4 Additional analyses of Chain-of-Frames
-Figure 6: Diversity of training data.
-Combining CoF examples extracted from real and synthetic videos provides better results on almost all benchmarks compared to using a single source.
-For all models we use 164k training samples, and fine-tune an InternVL2.5-4B model.
-Influence of training data.
-The training data of CoF-Data combines chain-of-frames from real and synthetic videos (total of 164k samples).
-To evaluate the impact of dataset diversity, we construct datasets of the same size (164k samples) but from a single source, i.e., either CoF-Datareal or CoF-Datasynth, and fine-tune InternVL2.5-4B on them.
-We compare these two models to standard CoF-InternVL2.5-4B (trained on CoF-Data) in Fig.
-6.
-Using the combined dataset outperforms single-source datasets on all benchmarks except for EventHallusion, demonstrating the importance of diversity in the reasoning traces used for training.
-Between the two individual datasets, the model trained on synthetic videos outperforms the one trained on real videos in three out of five cases, while being worse in just one.
-This suggests that further improvements might come from expanding the tasks covered by the synthetic data, which can even be generated at low cost.
-CoF reasoning at inference time.
+Finally, our proposed model, CoF-InternVL2.5-4B, fine-tuned on the CoF reasoning traces, achieves the highest accuracy across all benchmarks, with improvements ranging from 4.8% to 8.7% over the original InternVL2.5-4B.
+These results demonstrate that our explicit temporal ground approach, chain-of-frames, offers consistent benefits across diverse tasks and domains and can serve as a strong alternative to key-frame-retrieval methods.
+5.3 CoF reasoning at inference time
 To better understand how the CoF models utilize frame references during inference, we track the number of frames referenced in the reasoning trace per answer.
 In Fig.
-7, we report how many answers, generated by our CoF-InternVL2.5-4B and CoF-InternVL3-8B models across all evaluation benchmarks, contain $N$ frames for $N=1,\ldots,10$.
-We can see that both models learn to produce, from a relatively small training set, reasoning traces which refer to frame IDs (see qualitative examples in App.
+7, we report how many answers, generated by our CoF-InternVL3-8B model across all evaluation benchmarks, contain $N$ frames for $N=1,\ldots,10$.
+We can see that our model learns to produce, from a relatively small training set, reasoning traces which refer to frame IDs (see qualitative examples in App.
 C).
-In particular, CoF-InternVL2.5-4B includes at least one frame reference in 68.7% of the answers, while CoF-InternVL3-8B does it in 76.9% of the cases.
-This suggests that larger models might benefit more from fine-tuning with frame-grounded reasoning: in fact, the CoF-InternVL3-8B model shows both higher use of frame references and larger gains in accuracy (see Fig.
-1).
-Finally, the models can use frame references selectively rather than uniformly across tasks ( distribution by benchmark in App.
+Moreover, the model can use frame references selectively rather than uniformly across tasks (distribution by benchmark in App.
 B), depending on the type of questions.
-Figure 7: CoF reasoning at inference time.
-For both CoF-InternVL2.5-4B (left plot) and CoF-InternVL3-8B (right), we show the distribution of the number of frame references generated at inference time: both models learn to produce chain-of-frames during evaluation.
+Figure 7: CoF reasoning at inference time. We show the distribution of the number of frame references generated by CoF-InternVL3-8B at inference time.
 6 Conclusion
-We have introduced chain-on-frames (CoF), a new approach to encourage video LLMs to produce temporally grounded reasoning before providing the final answer.
-Compared to existing works, CoF does not require complex ad-hoc inference frameworks or auxiliary models, and we show that its training data can be efficiently extracted from both real and synthetic videos.
+We have introduced chain-of-frames (CoF), a new approach to encourage video LLMs to produce temporally grounded reasoning before providing the final answer.
+Compared to existing works, CoF does not require complex ad-hoc inference frameworks or auxiliary models, and we show that its training data can be extracted more efficiently and accurately from both real and synthetic videos.
 Our models fine-tuned on CoF-Data outperform across multiple benchmarks those obtained with alternative methods for reasoning, and even achieve results better than or similar to leading video LLMs.
@@
 Limitations.
-We have currently applied CoF on the InternVL models since their video encoding is well-suited for our reasoning with frame references, and they achieve state-of-the-art performance.
-It is an interesting open question how to customize our approach to other types of video LLMs.
+Our CoF training data is suitable for finetuning video LLMs that operate with a fixed FPS, ensuring that the frame indices referenced in reasoning traces align with the video frames provided to the language model.
+Consequently, an open question is how to adapt our method to video LLMs—such as Qwen2.5-VL [4]—that employ dynamic frame rates, while preprocessing the video.
 Acknowledgements
@@
 CoF from real videos (CoF-Datareal).
-To generate question-reasoning-answer triplets, we prompt Llama-3.1-8B-Instruct (Meta AI, 2024) using the instruction shown in Fig.
+To generate question-reasoning-answer triplets, we prompt Llama-3.1-8B-Instruct [27] using the instruction shown in Fig.
 9 along with frame-aware video captions from the VideoEspresso dataset (see Fig.
@@
 Two examples from CoF-Datareal are shown in Fig.
-12.
+13.
 CoF from synthetic videos (CoF-Datasynth).
@@
 The questions, answers and reasoning traces are generated with the manually written templates shown in Fig.
-9, making the data collection process particularly simple and fast.
+10, making the data collection process particularly simple and fast.
 Examples from CoF-Datasynth are shown in Fig.
-13.
+14.
+Figure 8: Distribution of frame references in the Chain-of-Frames training data.
+The left pie chart illustrates the distribution for CoF-Datareal, having fewer frames per reasoning trace, whereas CoF-DATA-synth demonstrates a more balanced frame distribution due to controlled synthetic video generation.
+The right pie chart shows the overall distribution for the CoF-Data.
+Final dataset (CoF-Data).
+Our training data, CoF-Data, has a total of 164,186 samples, comprising 103,683 samples from the CoF-Datareal dataset, which is based on real-world videos, and 60,503 samples from the CoF-Datasynth dataset of synthetic videos.
+Fig.
+8 shows the distribution of how many frames are referenced in the reasoning traces, both for the final dataset and the individual splits.
+The CoF-Datasynth exhibits a more balanced distribution compared to the automatically generated COF-CoF-Datareal: this highlights that using synthetic videos allows us to better control various aspects of the data.
 Prompt
@@
 Video-MME.
-Video-MME (Fu et al., 2024) offers a diverse range of video types, covering six primary visual domains and 30 subfields to support broad scenario generalizability.
+Video-MME [10] offers a diverse range of video types, covering six primary visual domains and 30 subfields to support broad scenario generalizability.
 It also introduces variation in temporal length, including short (under 2 minutes), medium (4-15 minutes), and long (30-60 minutes) videos.
@@
 Li et al.
-(2023b) presents a comprehensive benchmark for multimodal video understanding, encompassing 20 challenging tasks that require more than single-frame analysis.
+[26] presents a comprehensive benchmark for multimodal video understanding, encompassing 20 challenging tasks that require more than single-frame analysis.
 It is specifically designed to evaluate a model’s ability to understand temporal dynamics across video sequences.
 VSI-Bench.
-This benchmark (Yang et al., 2024) is designed to quantitatively assess the visual-spatial intelligence of multimodal large language models.
+This benchmark [38] is designed to quantitatively assess the visual-spatial intelligence of multimodal large language models.
 Built from over 5,000 high-quality question-answer pairs across 288 real-world indoor videos, VSI-Bench spans diverse environments such as homes, offices, and industrial spaces.
@@
 Out of these tasks included in this benchmark, relative distance, appearance order, relative directory, and route planning come with multiple-choice questions while the other four require an open-ended quantitative answer.
-To better evaluate the proximity of the model’s prediction with the correct answer, Yang et al.
-(2024) proposes using mean relative accuracy ($\mathcal{MRA}$).
+To better evaluate the proximity of the model’s prediction with the correct answer, [38] proposes using mean relative accuracy ($\mathcal{MRA}$).
 Given a model’s prediction $\hat{y}$ and ground truth $y$, relative accuracy is calculated by:
@@
 VidHal.
-To evaluate video-based hallucinations in video LLMs, we use VidHal (Choong et al., 2024), a multiple-choice benchmark that features video instances drawn from public video understanding datasets, covering a diverse array of temporal concepts and aspects such as entity actions and event sequences.
+To evaluate video-based hallucinations in video LLMs, we use VidHal [8], a multiple-choice benchmark that features video instances drawn from public video understanding datasets, covering a diverse array of temporal concepts and aspects such as entity actions and event sequences.
 EventHallusion.
 Zhang et al.
-(2024a) introduce EventHallusion, from which we use the binary-choice questions designed to systematically assess event-related hallucinations in video LLMs.
+[40] introduce EventHallusion, from which we use the binary-choice questions designed to systematically assess event-related hallucinations in video LLMs.
 From a hallucination attribution standpoint, it is specifically curated to evaluate a model’s susceptibility to language priors and vision-language correlation biases.
@@
 A.3 Chain-of-Frames model
+CoF-InternVL.
 For InternVL2.5-4B, we fully fine-tune both the LLM and the projection modules, keeping the vision encoder frozen.
-In contrast, for InternVL3-8B, we adopt LoRA-based fine-tuning (Hu et al., 2022) to reduce memory consumption.
+In contrast, for InternVL3-8B, we adopt LoRA-based fine-tuning [15] to reduce memory consumption.
 All other training configurations remain consistent across both models.
 Training is conducted on a single H100 node equipped with 4 GPUs, using a learning rate of $2\times 10^{-6}$, a batch size of 2, and a single epoch.
+CoF-Phi-3.5-Vision-4B.
+To test the generalizability of our approach beyond the InternVL family, we employ Phi-3.5-Vision-4B [1], a mobile-scale multimodal LLM that demonstrates strong performance in language reasoning, as a third baseline model.
+Phi-3.5-Vision-4B uses distinct, indexed image placeholder tokens such as <imagei-i>, instead of repeating a generic <image> token, because the model must uniquely align each image embedding with a specific position in the text sequence.
+Each placeholder is a separate token in the tokenizer, allowing the model to reliably map image embeddings to their corresponding locations and to correctly understand references to the images.
+Phi-3.5-Vision’s capability to understand references to the images makes it a good candidate for our base model.
+Similar to InternVL2.5-4B, we fully fine-tune both the LLM and the projection modules, keeping the vision encoder frozen.
+The other training configurations remain consistent across all baselines.
 Appendix B Additional Experiments
-Comparison to additional baselines.
-We could not include the original models from other works (Wang et al., 2024b; Han et al., 2024; Fei et al., 2024; Hu et al., 2025b) in the main evaluation in Table 2 because they are not accessible to the public, and they do not report results on the five benchmarks we use.
-However, we provide comparative analyses to M-LLM (Hu et al., 2025b) and Video-of-Thought (Fei et al., 2024) on Video-MME and NextQA (Xiao et al., 2021) using the results they report.
-In Table 3, we see that our CoF-based models, including CoF-InternVL2.5-4B which has fewer parameters than the competitors, outperform these baselines on both benchmarks.
-Moreover, compared to the most recent model, M-LLM, our method achieves a greater improvement (4.9% vs 0.8%) on NextQA despite starting from a stronger base model.
-Table 3: Comparison to additional baselines.
-Since the models from Hu et al.
-(2025b) and Fei et al.
-(2024) are not publicly available, we compare the results they report on Video-MME and NextQA to those of our CoF models.
-We also report the video LLM each model is based on and its performance, as well as the improvement of each approach over the original model.
-Both our CoF models outperform the existing approaches on the two benchmarks.
-Backbone Model Video-MME NextQA
-Video-LLaVA-7B Original 39.9 66.3
-Video-of-Thought (Fei et al., 2024) - 76.0  9.7$\uparrow$
-Qwen2-VL-7B Original 58.1 77.6
-M-LLM (Hu et al., 2025b) 58.7  0.6$\uparrow$ 78.4  0.8$\uparrow$
-InternVL2.5-4B Original 54.9 75.3
-SFT with CoF (ours) 59.7  4.8$\uparrow$ 79.6  4.3$\uparrow$
-InternVL3-8B Original 66.5 82.4
-SFT with CoF (ours) 75.38.8$\uparrow$ 87.34.9$\uparrow$
+CoF-Phi-3.5-Vision-4B results.
+Tab.
+4 compares the baseline Phi-3.5-Vision-4B model with its CoF-enhanced variant.
+Across all benchmarks, CoF-Phi-3.5-Vision-4B consistently surpasses the corresponding base model.
+The largest improvements appear on the VSI-Bench benchmark, where CoF-Phi-3.5-Vision-4B achieves better performance compared to LLaVA-OneVision-7B and Qwen2-VL-7B (Tab.
+1) despite using substantially fewer parameters.
+These results demonstrate that our chain-of-frames approach generalizes effectively across architectures and model families.
 Effect of CoT prompting.
-An extended version of Table 1 is presented in Table 4.
-For all baselines, we report results using two prompting strategies, either standard (indicated by $\bm{\star}$) or chain-of-thought (indicated by $\clubsuit$, the prompt is shown in Fig.
-10).
-For our SFT with CoF models, we always use CoT prompting.
-When considering InternVL2.5-4B, CoT prompting alone improves the accuracy of the original models on four out of five benchmarks compared to the original model.
-However, this improvement does not hold for the SFT with QA only variant: we hypothesize that fine-tuning solely on QA data negatively impacts the reasoning capabilities of the baseline model.
-On the other hand, incorporating reasoning traces into the training data (SFT with CoT) generally enhances the model’s reasoning capabilities, and using CoT prompting is beneficial except for the EventHallusion benchmark.
-CoT prompting improves the results also for the original InternVL3-8B on most benchmarks.
-Finally, our models (SFT with CoF) outperform the baseline across all benchmarks.
-Table 4: Effect of CoT prompting.
-For all baselines, we report results using two prompting strategies, i.e.
-standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix CoT prompting for our CoF models.
+[…]
 Model Prompt VSI-Bench Video-MME MVBench VidHal EventHall
 InternVL2.5-4B
-Original $\bm{\star}$ 31.8 54.9 70.8 74.0 62.5
+[…]
 $\clubsuit$ 33.5 54.7 71.5 77.0 67.4
-SFT with QA only $\bm{\star}$ 31.8 55.4 70.3 73.6 63.1
+[…]
 $\clubsuit$ 31.8 54.5 73.4 64.1 57.7
-SFT with CoT $\bm{\star}$ 31.1 52.6 69.6 74.4 62.5
+[…]
 $\clubsuit$ 34.3 58.6 73.7 77.9 53.1
@@
 InternVL3-8B
-Original $\bm{\star}$ 41.0 62.3 72.0 80.9 72.1
+[…]
 $\clubsuit$ 40.2 66.5 74.3 61.6 73.9
-SFT with CoF (ours) $\clubsuit$ 51.3 75.3 90.4 79.5 78.7
+SFT with CoF (ours) $\clubsuit$ 51.3 73.7 77.1 79.5 78.7
+Phi-3.5-Vision-4B
+[…]
+$\clubsuit$ 26.4 39.0 46.6 43.1 51.3
+SFT with CoF (ours) $\clubsuit$ 32.8 52.6 51.8 55.5 59.2
+[…]
 Detailed results over benchmark splits.
-For completeness, we report the fine-grained results over the various splits of VSI-Bench (Table 5), Video-MME (Table 6), MVBench (Table 7), and EventHallusion (Table 8).
-Moreover, for the baseline models, we report results using two prompting strategies, i.e., standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$).
+[…]
 Detailed results of CoF reasoning at inference time.
 In Fig.
-11, we show statistics of how many frames are referenced in the reasoning traces generated by CoF-InternVL3-8B.
+12, we show statistics of how many frames are referenced in the reasoning traces generated by CoF-InternVL3-8B.
 To complement Fig.
@@
 Figure 12: CoF reasoning at inference time. For each benchmark, we show the frequency of the number of frames referenced in the reasoning traces of CoF-8B.
-Table 5: Detailed results on the VSI-Bench benchmark.
-For the baselines, we report results using two prompting strategies, i.e.
-standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
 Model Prompt Obj. Count Abs. Dist. Obj. Size Room Size Rel. Dist. App. Order Rel. Dir. Route Plan Avg
 InternVL2.5-4B
-Original $\bm{\star}$ 29.2 31.2 45.5 20.4 35.4 23.2 41.4 27.8 31.8
+[…]
 $\clubsuit$ 36.0 17.1 38.1 29.8 34.2 30.3 52.2 29.9 33.5
-SFT with QA only $\bm{\star}$ 22.7 30.7 44.0 26.0 36.3 22.2 39.4 33.0 31.8
+[…]
 $\clubsuit$ 34.9 18.6 38.8 23.2 37.0 28.2 47.1 26.3 31.8
-SFT with CoT $\bm{\star}$ 31.5 22.0 41.6 27.9 36.8 21.2 40.4 27.3 31.1
+[…]
 $\clubsuit$ 39.1 19.5 36.1 26.9 36.1 30.1 57.1 29.4 34.3
@@
 InternVL3-8B
-Original $\bm{\star}$ 58.6 28.5 49.5 43.3 47.0 38.5 31.3 31.4 41.0
+[…]
 $\clubsuit$ 55.2 29.5 38.1 32.6 42.8 47.1 47.5 28.4 40.2
 SFT with CoF (ours) $\clubsuit$ 61.8 34.2 37.7 26.7 66.8 43.4 83.9 55.7 51.3
-Table 6: Detailed results on the Video-MME benchmark.
-For the baselines, we report results using two prompting strategies, i.e.
-standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
+Phi-3.5-Vision-4B
+[…]
+$\clubsuit$ 27.3 24.1 32.6 22.6 20.5 20.9 35.2 28.3 26.4
+SFT with CoF (ours) $\clubsuit$ 31.5 27.9 37.9 26.3 33.5 22.7 47.3 35.1 32.8
+[…]
 Model Prompt Short (900) Medium (900) Long (900) Avg
 InternVL2.5-4B
-Original $\bm{\star}$ 64.9 52.7 47.2 54.9
+[…]
 $\clubsuit$ 64.0 53.2 47.0 54.7
-SFT with QA only $\bm{\star}$ 68.0 53.6 44.8 55.5
+[…]
 $\clubsuit$ 66.8 53.1 43.6 54.5
-SFT with CoT $\bm{\star}$ 64.3 51.8 41.8 52.6
+[…]
 $\clubsuit$ 70.4 55.7 49.6 58.6
@@
 InternVL3-8B
-Original $\bm{\star}$ 73.0 61.7 52.1 62.3
+[…]
 $\clubsuit$ 75.3 65.3 59.0 66.6
-SFT with CoF (ours) $\clubsuit$ 80.9 74.4 70.4 75.3
-Table 7: Detailed results on the MVBench benchmark.
-For the baselines, we report results using two prompting strategies, i.e.
-standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
+SFT with CoF (ours) $\clubsuit$ 79.3 71.7 70.0 73.7
+Phi-3.5-Vision-4B
+[…]
+$\clubsuit$ 46.0 34.9 36.1 39.0
+SFT with CoF (ours) $\clubsuit$ 63.1 51.9 42.7 52.6
+[…]
 Model Prompt AA AC AL AP AS CO CI EN FA MA
 InternVL2.5-4B
-Original $\bm{\star}$ 89.5 54.0 44.0 75.0 82.9 62.5 79.0 29.0 46.0 97.5
+[…]
 $\clubsuit$ 88.5 50.0 46.5 77.0 81.4 67.0 78.0 34.5 60.5 99.0
-SFT with QA only $\bm{\star}$ 90.0 55.0 44.0 76.5 81.9 63.5 75.0 33.0 46.0 98.5
+[…]
 $\clubsuit$ 90.5 50.0 48.5 78.5 84.0 67.5 80.0 38.5 71.5 99.5
-SFT with CoT $\bm{\star}$ 87.0 53.0 35.5 76.5 81.4 62.0 76.5 31.5 43.5 98.0
+[…]
 $\clubsuit$ 90.5 46.5 57.5 85.0 83.5 67.0 80.5 38.5 73.5 98.5
@@
 InternVL3-8B
-Original $\bm{\star}$ 90.0 42.0 44.5 83.0 82.5 75.5 78.0 38.5 45.0 98.0
-$\clubsuit$ 77.5 45.0 42.5 87.5 97.9 80.5 89.0 33.5 54.0 99.0
-SFT with CoF (ours) $\clubsuit$ 96.5 95.5 58.0 99.5 100.0 99.0 77.5 79.5 87.5 96.5
+[…]
+$\clubsuit$ 77.5 45.0 42.5 87.5 85.1 80.5 89.0 33.5 47.0 99.0
+SFT with CoF (ours) $\clubsuit$ 96.5 50.5 49.5 89.5 91.0 91.5 77.5 45.5 59.5 96.5
+Phi-3.5-Vision-4B
+[…]
+$\clubsuit$ 61.0 49.5 27.5 43.0 37.8 42.0 35.0 36.5 36.5 61.5
+SFT with CoF (ours) $\clubsuit$ 68.0 51.5 38.5 49.5 42.02 49.5 39.5 44.0 38.5 68.0
 Model Prompt MC MD OE OI OS ST SC UA Avg
 InternVL2.5-4B
-Original $\bm{\star}$ 88.5 73.0 96.5 83.5 39.5 92.0 57.5 85.0 70.8
+[…]
 $\clubsuit$ 86.5 72.5 96.0 81.5 40.5 91.5 58.0 78.0 71.5
-SFT with QA only $\bm{\star}$ 87.5 75.0 96.5 82.5 41.0 91.5 59.5 85.5 70.3
+[…]
 $\clubsuit$ 86.5 75.5 96.5 86.5 42.0 92.0 52.5 82.0 73.4
-SFT with CoT $\bm{\star}$ 89.0 72.5 96.5 82.0 38.0 91.5 56.5 82.0 69.6
+[…]
 $\clubsuit$ 86.5 72.0 95.5 86.5 40.5 92.0 52.5 80.5 73.7
@@
 InternVL3-8B
-Original $\bm{\star}$ 60.5 89.0 97.0 85.5 39.5 92.5 69.0 81.0 72.0
-$\clubsuit$ 70.5 91.0 97.5 87.5 41.5 92.5 77.0 73.5 74.3
-SFT with CoF (ours) $\clubsuit$ 85.0 89.0 90.4 96.5 95.5 94.0 99.0 88.0 90.4
-Table 8: Detailed results on the VidHal and EventHallusion benchmarks.
-For the baselines, we report results using two prompting strategies, i.e.
-standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
+[…]
+$\clubsuit$ 63.5 89.0 97.47 87.5 41.5 92.5 73.0 73.5 72.5
+SFT with CoF (ours) $\clubsuit$ 67.0 91.0 90.4 90.5 45.5 94.0 78.0 84.0 77.1
+Phi-3.5-Vision-4B
+[…]
+$\clubsuit$ 35.5 38.5 53.0 43.5 28.5 89.0 47.5 72.5 46.6
+SFT with CoF (ours) $\clubsuit$ 36.5 45.5 61.1 46.5 33.5 90.0 53.3 76.5 51.8
+[…]
 Model Prompt VidHal EventHallusion
@@
 InternVL2.5-4B
-Original $\bm{\star}$ 74.0 48.3 91.2 48.2 62.5
+[…]
 $\clubsuit$ 77.0 44.7 80.4 77.2 67.4
-SFT with QA only $\bm{\star}$ 73.6 48.3 89.2 51.8 63.1
+[…]
 $\clubsuit$ 64.1 47.4 75.5 50.3 57.7
-SFT with CoT $\bm{\star}$ 74.4 49.1 91.2 47.1 62.5
+[…]
 $\clubsuit$ 77.9 39.5 71.6 48.2 53.1
@@
 InternVL3-8B
-Original $\bm{\star}$ 80.9 52.6 91.2 72.5 72.1
+[…]
 $\clubsuit$ 61.6 57.0 94.1 70.5 73.9
 SFT with CoF (ours) $\clubsuit$ 79.5 57.9 92.2 86.0 78.7
+Phi-3.5-Vision-4B
+[…]
+$\clubsuit$ 43.1 22.8 64.7 66.3 51.3
+SFT with CoF (ours) $\clubsuit$ 55.5 34.2 73.5 70.0 59.2
+[…]
 Appendix C Additional Figures
@@
 15 and samples from hallucination benchmarks in Fig.
-15.
+16.
 Figure 13: CoF training data. Examples of the training data from CoF-Datareal, including relevant frames, questions, answers, and reasoning traces.